### v1

# Gauge-independent gravitational waves from a minimal dark $U(1)$ sector with viable dark matter candidates

Abstract
Searches for stochastic gravitational wave backgrounds from first-order phase transitions offer a powerful probe of hidden sectors, but quantitative predictions in gauge theories are obstructed by the gauge dependence of the finite-temperature effective potential and the associated tunneling action. We study a minimal gauged $U(1)$ dark sector containing a dark Higgs and a dark photon, optionally supplemented by a vectorlike dark fermion, coupled to the Standard Model through Higgs portal or kinetic mixing. Using the Nielsen identity together with a controlled derivative expansion and power counting, we construct a gauge-independent effective action in the high- and low-temperature limits, enabling model-intrinsic nucleation dynamics and robust gravitational wave predictions. We perform dedicated Monte Carlo scans in both limits and map viable microscopic parameters to detector-facing peak frequencies and amplitudes, spanning bands relevant to pulsar timing arrays and planned space-based interferometers. In our scans, supercooled transitions typically produce the strongest signals, whereas parametrically high-temperature transitions are comparatively rare and tend to be weak. We further connect the phase transition phenomenology to viable dark matter candidates within the same minimal field content, providing benchmark targets for dark photon dark matter and dark fermion dark matter, and highlighting their complementarity with gravitational wave observables. Overall, our results provide an end-to-end, gauge-independent pipeline from a minimal hidden sector Lagrangian to gravitational wave spectra and cosmologically viable dark matter benchmarks, yielding the most reliable and concrete predictions to date for a minimal gauged $U(1)$ dark sector.
1 Introduction
The discovery of gravitational waves provides a novel probe of hidden sectors in which dark matter may reside. Searches for stochastic gravitational wave backgrounds produced by first-order phase transitions, either from the Standard Model (SM) augmented by new physics [1, 2, 3, 4, 5, 6, 7], or in a dark sector [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 33, 31, 32, 34, 35, 36], offer a powerful and complementary pathway to uncovering new physics, especially in light of the continuing non-detection of dark matter in conventional experiments. Gravitational wave observations thus constitute a key component of a multi-messenger program to test particle physics in the early Universe. Accordingly, the interplay between gravitational wave phenomenology and physics beyond the SM, particularly in scenarios aimed at explaining the nature of dark matter, has become one of the most active frontiers of current research.
In well-motivated dark sector models, gravitational wave from a first-order phase transition is particularly compelling: New gauge symmetries and scalar fields are ubiquitous in beyond the SM frameworks, and their symmetry breaking can naturally occur at scales both below and above the electroweak scale, leading to gravitational wave signals that span a wide range of frequencies. Consequently, it is essential to establish robust and concrete predictions for gravitational wave signals from dark sectors that can be confronted with data across multiple frequency bands. If detected, such a signal can reveal information about the underlying dark sector parameters including the gauge coupling, the scalar quartic coupling, and the scalar vacuum expectation value (vev), that may be difficult or impossible to access directly through collider probes or conventional dark matter searches.
However, a central obstacle to making reliable gravitational wave predictions in gauge theories is the well-known gauge dependence of the finite-temperature effective potential and the associated tunneling action. The Higgs effective potential as a function of an arbitrary background field is not a physical observable. In $R_{\xi}$ gauges, the result depends explicitly on the gauge-fixing parameter $\xi$, and thus a conventional computation based directly on the gauge-dependent effective potential leads to gauge-dependent gravitational wave spectra that can hardly be regarded as genuine “predictions”. This gauge ambiguity obscures the interpretation of parameter scans and hampers the extraction of robust, model-intrinsic conclusions.
To address this issue, we adopt a gauge-independent formulation of the effective action based on the Nielsen identity [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47], with particular emphasis on the low-temperature treatment developed in [34]. The key payoff is a well-defined, gauge-independent tunneling action and thus gauge-independent predictions for the phase transition and the resulting gravitational wave signal within a controlled power counting scheme. We provide a detailed discussion of the validity of the derivative expansion of the effective action, along with the associated gauge coupling power counting rules, and the scaling relations between the gauge coupling and the quartic coupling at zero, high, and low temperatures, thereby clarifying several common sources of confusion in the literature. Building on the gauge-independent effective action in the relevant high- and low-temperature limits, we perform dedicated Monte Carlo scans in both regimes. This yields concrete, detector-facing predictions: each viable parameter point maps to a peak frequency and amplitude of the stochastic background, enabling a direct comparison with the sensitivity targets of current and planned experiments across various frequency bands.
We focus in particular on supercooled first-order phase transitions, which are among the strongest and most distinctive sources of a stochastic gravitational wave background [48, 49, 50, 51, 52, 53, 54, 19, 20, 55, 56, 57, 58, 59, 21, 60, 61, 62, 63, 33, 34, 35, 64]. In such transitions, the nucleation and percolation temperatures satisfy $T_{n,p}\ll T_{c}$, which typically enhances the gravitational wave amplitude and shifts the peak frequency to lower values, often bringing the signal into the nanohertz band and thereby attracting substantial recent attention. Moreover, models that undergo strong supercooling often arise in nearly flat potentials, for which the tunneling action remains large and nucleation is delayed. In many such scenarios, the portal couplings to the SM are also small, making these models challenging to probe with collider searches or conventional dark matter experiments. Gravitational wave observations therefore become a leading discovery channel and can provide compelling evidence for this region of parameter space.
In this work we focus on the simplest representative setup: a minimal $U(1)_{x}$ dark sector featuring a dark Higgs field and a dark photon, and optionally a dark fermion, together with portal interactions connecting the hidden sector to the SM. This minimality is not merely aesthetic. It makes the parameter dependence transparent, allows for a systematic exploration of distinct thermal regimes, and admits dark matter candidates whose cosmological evolution can be computed consistently alongside the phase transition dynamics. In particular, we highlight two viable dark matter candidates already present in the minimal field content: (1) dark photon dark matter, which requires vanishing or negligibly small kinetic mixing; and (2) dark fermion dark matter, which relies on a non-negligible dark photon portal.
This paper therefore provides robust and predictive results for gravitational waves generated by the symmetry breaking of a minimal gauged $U(1)$ dark sector with a viable single dark matter candidate. Such results are significant in the current high energy physics research, where no compelling hints of new physics have emerged from any experiments, and precision tests as well as multi-messenger searches are becoming increasingly essential for new physics discovery.
The remainder of this paper is organized as follows. In Section 2 we introduce the minimal dark $U(1)_{x}$ sector and portal interactions that connect it to SM states. In Section 3 we review the gauge dependence of the finite-temperature effective potential and the construction of gauge-independent effective actions based on the Nielsen identity, with particular emphasis on the low-temperature regime relevant for supercooled phase transitions. In Section 4 we present the computation of gravitational wave signatures from first-order phase transitions, including the tunneling action and transition rate, characteristic temperatures, hydrodynamic parameters, and the resulting gravitational wave power spectra, and we compare gauge-dependent result and gauge-independent predictions. In Section 5 we analyze the cosmological evolution of dark sector particles, derive the coupled Boltzmann equations for the dark matter scenarios of interest, and present benchmark models and phenomenological implications. Finally, we summarize our conclusions and present an outlook in Section 6.
2 The $U(1)_{x}$ extension and portals to the SM
In this section, we review the minimal $U(1)_{x}$ extension of the SM, featuring a dark Higgs, a dark photon, and optionally dark fermions, with their portal interactions to SM particles.
2.1 The $U(1)_{x}$ extension of the SM
We consider a $U(1)_{x}$ gauge extension of the SM. The full Lagrangian can be written as
$\mathcal{L}=\mathcal{L}_{{\rm SM}}+\mathcal{L}_{{\rm hid}}+\mathcal{L}_{{\rm mix}}\,$, (2.1)
where $\mathcal{L}_{{\rm SM}}$ and $\mathcal{L}_{{\rm hid}}$ is given by
$\mathcal{L}_{{\rm SM}}\supset$ $-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\mu_{{\rm SM}}^{2}H^{\dagger}H+\lambda_{{\rm SM}}(H^{\dagger}H)^{2}\,$, (2.2)
$\mathcal{L}_{\mathrm{hid}}=$ $\;-\frac{1}{4}F_{x\,\mu\nu}F_{x}^{\mu\nu}+\bigl|(\partial_{\mu}-ig_{x}A_{\mu}^{\prime})\Phi\bigr|^{2}+\mu_{x}^{2}\Phi^{\ast}\Phi-\lambda_{x}(\Phi^{\ast}\Phi)^{2}$
$+g_{x}\overline{\chi}\gamma^{\mu}\chi A_{\mu}^{\prime}+\overline{\chi}(i\gamma^{\mu}\partial_{\mu}-m_{\chi})\chi\,$, (2.3)
where $F_{\mu\nu},F_{x\,\mu\nu}$ are the field strengths of the hypercharge and $U(1)_{x}$ gauge fields respectively, $H$ is the SM Higgs doublet and $\Phi$ is the $U(1)_{x}$ Higgs field with $U(1)_{x}$ charge +1+1. $A_{\mu}^{\prime}$ is the $U(1)_{x}$ gauge field, and $\chi$ is a dark fermion with vector mass $m_{\chi}$ and $U(1)_{x}$ charge $Q_{x}=+1$. The most important parameters relevant for the study of both $U(1)_{x}$ gravitational wave physics and dark matter physics are: the $U(1)_{x}$ gauge coupling $g_{x}$, the $U(1)_{x}$ Higgs vev $v_{x}$ and its quartic coupling $\lambda_{x}$.
The mixing terms contain kinetic mixing and Higgs mixing contributions, written as
$\mathcal{L}_{{\rm mix}}=\;-\frac{\delta}{2}F_{\mu\nu}F_{x}^{\mu\nu}-\lambda_{{\rm mix}}(H^{\dagger}H)(\Phi^{\ast}\Phi)^{2}\,$, (2.4)
where $\delta$ and $\lambda_{{\rm mix}}$ are mixing parameters, both receiving stringent constraints from experiments. The experimental constraint of $\lambda_{{\rm mix}}$ will be discussed in Section 5.2. The kinetic mixing can result in from either integrating out bifundamental fields charged under both the $U(1)_{Y}$ and $U(1)_{x}$, which can be absent; or from graviton loop effects, which are in general suppressed. We focus on two classes of dark matter candidates in this paper:
  • Dark photon dark matter, which requires vanishing or extremely small kinetic mixing, $\delta\lesssim 10^{-16}$.
  • Dark fermion dark matter, which requires a non-negligible kinetic mixing.
Expanding $\Phi$ around the background field $\phi_{c}$ as
$\Phi=\frac{1}{\sqrt{2}}\bigl(\phi_{c}+h_{x}+iG\bigr)\,$, (2.5)
where $h_{x}$ and $G$ are the dark Higgs and Goldstone boson respectively, we obtain
$\bigl|(\partial_{\mu}-ig_{x}A_{\mu}^{\prime})\Phi\bigr|^{2}$ $=\frac{1}{2}(\partial_{\mu}h_{x}+g_{x}A_{\mu}^{\prime}G)^{2}+\frac{1}{2}\bigl[\partial_{\mu}G-g_{x}A_{\mu}^{\prime}(\phi_{x}+h_{x})\bigr]^{2}$
$=\frac{1}{2}(\partial_{\mu}h_{x})^{2}+\frac{1}{2}(\partial_{\mu}G)^{2}-g_{x}\phi_{c}(\partial_{\mu}G)A^{\prime\mu}+\cdots\,.$ (2.6)
Together with the gauge fixing term and the ghost terms,
$\mathcal{L}_{{\rm gf}}$ $=-\frac{1}{2\xi}\bigl(\partial_{\mu}A^{\prime}_{\mu}+\xi g_{x}\phi_{c}G\bigr)^{2}\,$, (2.7)
$\mathcal{L}_{{\rm ghost}}$ $=\overline{c}\,\bigl[-\partial^{2}-\xi g_{x}^{2}\phi_{c}(\phi_{c}+h_{x})\bigl]\,c\,$, (2.8)
where $\xi$ is the gauge parameter, $c(\overline{c})$ is the ghost (anti-ghost) field, the mixing between $A^{\prime}$ and $G$ in Eq. (2.6) is canceled by the cross term in $\mathcal{L}_{{\rm gf}}$. The field-dependent masses in the $R_{\xi}$ gauge are given by
$m_{A^{\prime}}^{2}$ $=g_{x}^{2}\phi_{c}^{2}\,,\qquad m_{A^{\prime}_{0}}^{2}=\xi m_{A^{\prime}}^{2}=\xi g_{x}^{2}\phi_{c}^{2}\,$, (2.9)
$m_{h_{x}}^{2}$ $=-\mu_{x}^{2}+3\lambda_{x}\phi_{c}^{2}\,$, (2.10)
$m_{c}^{2}$ $=\xi m_{A^{\prime}}^{2}=\xi g_{x}^{2}\phi_{c}^{2}\,$, (2.11)
$m_{G}^{2}$ $=-\mu_{x}^{2}+\lambda_{x}\phi_{c}^{2}+\xi m_{A^{\prime}}^{2}\,$, (2.12)
where the masses of the longitudinal (time-like) gauge component $A^{\prime}_{0}$, the ghost and the Goldstone boson depend on the gauge parameter $\xi$.
2.2 Higgs portal through scalar mixing
We now discuss the Higgs portal interactions from $H,\Phi$ quartic mixing term. The full scalar potential reads
$V(H,\phi_{x})=-\mu_{{\rm SM}}^{2}H^{\dagger}H-\mu_{x}^{2}\Phi^{\ast}\Phi+\lambda_{{\rm SM}}\bigl(H^{\dagger}H\bigr)^{2}+\lambda_{x}\bigl(\Phi^{\ast}\Phi\bigr)^{2}+\lambda_{{\rm mix}}\bigl(H^{\dagger}H\bigr)\bigl(\Phi^{\ast}\Phi\bigr)\,.$ (2.13)
Working in the unitary gauge, $H$ and $\Phi$ can be expanded around their minima as
$H=\frac{1}{\sqrt{2}}\left(\begin{array}[]{c}0\\ v_{{\rm SM}}+h^{0}\end{array}\right)\,,\qquad\Phi=\frac{1}{\sqrt{2}}\bigl(v_{x}+h_{x}^{0}\bigr)\,$, (2.14)
where $h^{0}$, $h_{x}^{0}$ are real scalars in the flavor eigenbasis. The vacuum expectation values $v_{{\rm SM}}$, $v_{x}$ are given by
$v_{{\rm SM}}^{2}$ $=\frac{4\mu_{{\rm SM}}^{2}\lambda_{x}-2\mu_{x}^{2}\lambda_{{\rm mix}}}{4\lambda_{x}\lambda_{{\rm SM}}-\lambda_{{\rm mix}}^{2}}\,$, (2.15)
$v_{x}^{2}$ $=\frac{4\mu_{x}^{2}\lambda_{{\rm SM}}-2\mu_{{\rm SM}}^{2}\lambda_{{\rm mix}}}{4\lambda_{x}\lambda_{{\rm SM}}-\lambda_{{\rm mix}}^{2}}\,.$ (2.16)
The kinetic and mass terms are then
$\bigl|(\partial_{\mu}-ig_{x}A_{\mu}^{\prime})\Phi\bigr|^{2}$ $=\frac{1}{2}\bigl(\partial_{\mu}h_{x}^{0}\bigr)^{2}+\frac{1}{2}g_{x}^{2}\Bigl[A^{\prime 2}\bigl(h_{x}^{0}\bigr)^{2}+2v_{x}A^{\prime 2}h_{x}^{0}+v_{x}^{2}A^{\prime 2}\Bigr]\,$, (2.17)
$-\mathcal{L}_{{\rm mass}}$ $=\frac{1}{2}\bigl(\begin{array}[]{cc}h^{0}&h_{x}^{0}\end{array}\bigr)\left(\begin{array}[]{cc}2\lambda_{{\rm SM}}v_{{\rm SM}}^{2}&\lambda_{{\rm mix}}v_{{\rm SM}}v_{x}\\ \lambda_{{\rm mix}}v_{{\rm SM}}v_{x}&2\lambda_{x}v_{x}^{2}\end{array}\right)\left(\begin{array}[]{c}h^{0}\\ h_{x}^{0}\end{array}\right)\,.$ (2.23)
An orthogonal transformation is performed to diagonalize the mass matrix, which transformed the original eigenbasis $(h^{0}\;\;h_{x}^{0})^{T}$ into mass eigenbasis $(h\;\;h_{x})^{T}$ as
$\left(\begin{array}[]{c}h\\ h_{x}\end{array}\right)=\left(\begin{array}[]{cc}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{array}\right)\left(\begin{array}[]{c}h^{0}\\ h_{x}^{0}\end{array}\right)\,$, (2.24)
where the mixing angle $\theta$ is calculated to be
$\theta=\frac{1}{2}\arctan\frac{\lambda_{{\rm mix}}v_{{\rm SM}}v_{x}}{\lambda_{x}v_{x}^{2}-\lambda_{{\rm SM}}v_{{\rm SM}}^{2}}\,.$ (2.25)
The mass eigenvalues are
$m_{h}^{2}$ $=\lambda_{{\rm SM}}v_{{\rm SM}}^{2}+\lambda_{x}v_{x}^{2}+(\lambda_{{\rm SM}}v_{{\rm SM}}^{2}-\lambda_{x}v_{x}^{2})\sqrt{1+\tan^{2}2\theta}\,$, (2.26)
$m_{h_{x}}^{2}$ $=\lambda_{{\rm SM}}v_{{\rm SM}}^{2}+\lambda_{x}v_{x}^{2}-(\lambda_{{\rm SM}}v_{{\rm SM}}^{2}-\lambda_{x}v_{x}^{2})\sqrt{1+\tan^{2}2\theta}\,.$ (2.27)
Define $\Delta m^{2}=m_{h_{x}}^{2}-m_{h}^{2}$ and we obtain
$v_{x}$ $=\frac{\Delta m^{2}\sin 2\theta}{2v_{{\rm SM}}\lambda_{{\rm mix}}}\,$, (2.28)
$\lambda_{{\rm SM}}$ $=\frac{m_{h}^{2}}{2v_{{\rm SM}}^{2}}+\frac{\Delta m^{2}\sin^{2}\theta}{2v_{{\rm SM}}^{2}}\,$, (2.29)
$\lambda_{x}$ $=\frac{2v_{{\rm SM}}^{2}\lambda_{{\rm mix}}^{2}}{\sin^{2}2\theta\,\Delta m^{2}}\biggl(\frac{m_{h_{x}}^{2}}{\Delta m^{2}}-\sin^{2}\theta\biggr)\,$, (2.30)
where the quartic coupling $\lambda_{{\rm mix}}$ must satisfy all current experimental constraints. In the mass eigenbasis, the scalar potential is now written as
$V(h,h_{x})=$ $\;C_{h^{3}}h^{3}+C_{h^{2}h_{x}}h^{2}h_{x}+C_{hh_{x}^{2}}hh_{x}^{2}+C_{h_{x}^{3}}h_{x}^{3}$
$+C_{h^{4}}h^{4}+C_{h^{3}h_{x}}h^{3}h_{x}+C_{h^{2}h_{x}^{2}}h^{2}h_{x}^{2}+C_{hh_{x}^{3}}hh_{x}^{3}+C_{h_{x}^{4}}h_{x}^{4}$
$+\frac{1}{2}\bigl(\begin{array}[]{cc}h&h_{x}\end{array}\bigr)\left(\begin{array}[]{cc}m_{h}^{2}&0\\ 0&m_{h_{x}}^{2}\end{array}\right)\left(\begin{array}[]{c}h\\ h_{x}\end{array}\right)\,$, (2.36)
where the trilinear couplings are given by
$C_{h^{3}}=\frac{m_{h}^{2}}{2v_{{\rm SM}}c_{\theta}}\Bigl(c_{\theta}^{4}-\frac{\lambda_{{\rm mix}}v_{{\rm SM}}^{2}}{\Delta m^{2}}s_{\theta}^{2}\Bigr)\,,\quad C_{h^{2}h_{x}}=\frac{2m_{h}^{2}+m_{h_{x}}^{2}}{2v_{{\rm SM}}}s_{\theta}\Bigl(c_{\theta}^{2}+\frac{\lambda_{{\rm mix}}v_{{\rm SM}}^{2}}{\Delta m^{2}}\Bigr)\,$, (2.37)
$C_{h_{x}^{3}}=\frac{m_{h_{x}}^{2}}{2v_{{\rm SM}}s_{\theta}}\Bigl(s_{\theta}^{4}+\frac{\lambda_{{\rm mix}}v_{{\rm SM}}^{2}}{\Delta m^{2}}c_{\theta}^{2}\Bigr)\,,\quad C_{hh_{x}^{2}}=\frac{m_{h}^{2}+2m_{h_{x}}^{2}}{2v_{{\rm SM}}}c_{\theta}\Bigl(s_{\theta}^{2}-\frac{\lambda_{{\rm mix}}v_{{\rm SM}}^{2}}{\Delta m^{2}}\Bigr)\,$, (2.38)
and the quartic couplings are
$C_{h^{4}}=\frac{1}{4}\Bigl(\frac{s_{2\theta}^{2}}{4}\lambda_{{\rm mix}}+c_{\theta}^{4}\lambda_{{\rm SM}}+s_{\theta}^{4}\lambda_{x}\Bigr)\,,\quad C_{h^{3}h_{x}}=\frac{1}{4}s_{2\theta}\bigl(-c_{2\theta}\lambda_{{\rm mix}}+2c_{\theta}^{2}\lambda_{{\rm SM}}-2s_{\theta}^{2}\lambda_{x}\bigr)\,$, (2.39)
$C_{h_{x}^{4}}=\frac{1}{4}\Bigl(\frac{s_{2\theta}^{2}}{4}\lambda_{{\rm mix}}+s_{\theta}^{4}\lambda_{{\rm SM}}+c_{\theta}^{4}\lambda_{x}\Bigr)\,,\quad C_{hh_{x}^{3}}=\frac{1}{4}s_{2\theta}\bigl(c_{2\theta}\lambda_{{\rm mix}}+2s_{\theta}^{2}\lambda_{{\rm SM}}-2c_{\theta}^{2}\lambda_{x}\bigr)\,$, (2.40)
$C_{h^{2}h_{x}^{2}}=\frac{1}{4}\Bigl[\bigl(c_{2\theta}^{2}+s_{2\theta}^{2}\bigr)\lambda_{{\rm mix}}+\frac{3}{2}s_{2\theta}^{2}\bigl(-\lambda_{{\rm mix}}+\lambda_{{\rm SM}}+\lambda_{x}\bigr)\Bigr]\,$, (2.41)
with $c_{\theta}\equiv\cos\theta$ and $s_{\theta}\equiv\sin\theta$. The $U(1)_{x}$ gauge boson $A^{\prime}$ acquires a mass $M_{A^{\prime}}=g_{x}v_{x}$, referred to as a dark photon $\gamma^{\prime}$. The couplings of $\gamma^{\prime}$ to two Higgs fields can be obtained from the kinetic term
$\bigl|(\partial_{\mu}-ig_{x}A_{\mu}^{\prime})\Phi\bigr|^{2}$ $=\frac{1}{2}\Bigl[s_{\theta}^{2}\bigl(\partial_{\mu}h\bigr)^{2}+c_{\theta}^{2}\bigl(\partial_{\mu}h_{x}\bigr)^{2}-2s_{\theta}c_{\theta}\bigl(\partial_{\mu}h\bigr)\bigl(\partial^{\mu}h_{x}\bigr)\Bigr]$
$\quad+\frac{1}{2}g_{x}^{2}A^{\prime 2}\Bigl[s_{\theta}^{2}h^{2}+c_{\theta}^{2}h_{x}^{2}-2s_{\theta}c_{\theta}h\,h_{x}+2v_{x}(s_{\theta}h+c_{\theta}h_{x})+v_{x}^{2}\Bigr]\,.$ (2.42)
The interactions of $h,h_{x}$ to SM particles are calculated to be
$\mathcal{L}=$ $\;\Bigl(M_{W}^{2}W^{\mu+}W_{\mu}^{-}+\frac{1}{2}M_{Z}^{2}Z^{\mu}Z_{\mu}\Bigr)\Bigl(1+\frac{h^{0}}{v_{{\rm SM}}}\Bigr)^{2}-\sum_{i}m_{i}\overline{f_{i}}f_{i}\frac{h^{0}}{v_{{\rm SM}}}$
$\to\;\Bigl(\frac{c_{\theta}^{2}}{v_{{\rm SM}}^{2}}h^{2}+\frac{2c_{\theta}}{v_{{\rm SM}}}h\Bigr)\Bigl(M_{W}^{2}W^{\mu+}W_{\mu}^{-}+\frac{1}{2}M_{Z}^{2}Z^{\mu}Z_{\mu}\Bigr)$
$\ \ +\Bigl(\frac{s_{\theta}^{2}}{v_{{\rm SM}}^{2}}h_{x}^{2}+\frac{2s_{\theta}}{v_{{\rm SM}}}h_{x}\Bigr)\Bigl(M_{W}^{2}W^{\mu+}W_{\mu}^{-}+\frac{1}{2}M_{Z}^{2}Z^{\mu}Z_{\mu}\Bigr)$
$\ \ +\frac{2c_{\theta}s_{\theta}}{v_{{\rm SM}}^{2}}hh_{x}\Bigl(M_{W}^{2}W^{\mu+}W_{\mu}^{-}+\frac{1}{2}M_{Z}^{2}Z^{\mu}Z_{\mu}\Bigr)-\frac{c_{\theta}h+s_{\theta}h_{x}}{v_{{\rm SM}}}\sum_{i}m_{i}\overline{f_{i}}f_{i}\,.$ (2.43)
2.3 $U(1)$ portal through kinetic mixing
We now discuss $U(1)_{x}$ portal interactions from the kinetic mixing. In the gauge eigenbasis $V^{T}=(C,B,A^{3})$ where $C,B$ and $A^{3}$ denote the dark $U(1)_{x}$ gauge field, SM hypercharge field, and the neutral $SU(2)_{L}$ gauge field respectively, the mixing matrices take the form [65, 66]
$\mathcal{K}=\left(\begin{array}[]{ccc}1&\delta&0\\ \delta&1&0\\ 0&0&1\end{array}\right)\,,\qquad M_{0}^{2}=\left(\begin{array}[]{ccc}g_{x}^{2}v_{x}^{2}&0&0\\ 0&\frac{1}{4}g_{Y}^{2}v_{{\rm SM}}^{2}&-\frac{1}{4}g_{2}g_{Y}v_{{\rm SM}}^{2}\\ 0&-\frac{1}{4}g_{2}g_{Y}v_{{\rm SM}}^{2}&\frac{1}{4}g_{2}^{2}v_{{\rm SM}}^{2}\end{array}\right)\,$, (2.44)
where $g_{2}$ and $g_{Y}$ denote the SM couplings. To diagonalize these two matrices simultaneously, we first perform a non-unitary transformation $K$ that diagonalizes the kinetic terms with $K^{T}\mathcal{K}K=\bm{1}$. Under the transformation $V^{\prime}=K^{-1}V$, the mass matrix becomes $M_{0}^{\prime 2}=K^{T}M_{0}^{2}K$. We then apply an orthogonal transformation $O$ such that $O^{T}K^{T}M_{0}^{2}KO$ is diagonal. The transformation matrices $K$ and $O$ are found to be
$K=\left(\begin{array}[]{ccc}c_{\delta}&0&0\\ -s_{\delta}&1&0\\ 0&0&1\end{array}\right)\,,\qquad O=\left(\begin{array}[]{ccc}1&0&0\\ 0&\cos\theta_{W}&-\sin\theta_{W}\\ 0&\sin\theta_{W}&\cos\theta_{W}\end{array}\right)\left(\begin{array}[]{ccc}\cos\psi&0&\sin\psi\\ 0&1&0\\ -\sin\psi&0&\cos\psi\end{array}\right)\,$, (2.45)
where $c_{\delta}=1/\sqrt{1-\delta^{2}}$, $s_{\delta}=\delta/\sqrt{1-\delta^{2}}$, $\cos\theta_{W}=g_{2}/\sqrt{g_{2}^{2}+g_{Y}^{2}}$, $\sin\theta_{W}=g_{Y}/\sqrt{g_{2}^{2}+g_{Y}^{2}}$. The angle $\psi$ is
$\psi=\frac{1}{2}\arctan\frac{2\delta\sqrt{1-\delta^{2}}\,\sin\theta_{W}}{1-\varepsilon^{2}-\delta^{2}(1+\sin^{2}\theta_{W})}\,$, (2.46)
where we define $\varepsilon^{2}\equiv M_{1}^{2}/M_{Z}^{2}$ with $M_{1}=g_{x}v_{x}$ and $M_{Z}=\tfrac{1}{2}v_{{\rm SM}}\sqrt{g_{2}^{2}+g_{Y}^{2}}$.
The full rotation matrix is thus
$\mathcal{R}\equiv KO=\left(\begin{array}[]{ccc}c_{\delta}\cos\psi&0&c_{\delta}\sin\psi\\ -s_{\delta}\cos\psi+\sin\psi\sin\theta_{W}&\cos\theta_{W}&-s_{\delta}\sin\psi-\cos\psi\sin\theta_{W}\\ -\sin\psi\cos\theta_{W}&\sin\theta_{W}&\cos\psi\cos\theta_{W}\end{array}\right)\,$, (2.47)
which diagonalizes the kinetic and mass matrices simultaneously and transform the original eigenbasis $V^{T}=(C,B,A^{3})$ to the mass eigenbasis $E^{T}=(A^{\prime},A_{\gamma},Z)$ with $E=\mathcal{R}^{T}V$. The resulting mass matrix is
$M^{2}=\mathcal{R}^{T}M_{0}^{2}\mathcal{R}=\left(\begin{array}[]{ccc}M_{A^{\prime}}^{2}&0&0\\ 0&0&0\\ 0&0&M_{Z}^{\prime 2}\end{array}\right)\,$, (2.48)
and the interactions of gauge bosons with fermions can be obtained from
$\mathcal{L}_{{\rm int}}=(g_{x}J_{x},\;g_{Y}J_{Y},\;g_{2}J_{3})\,V=(g_{x}J_{x},\;g_{Y}J_{Y},\;g_{2}J_{3})\,\mathcal{R}E\,.$ (2.49)
The neutral current couplings are summarized as follows
$\mathcal{L}_{\chi}$ $=g_{x}Q_{x}(\mathcal{R}_{11}A_{\mu}^{\prime}+\mathcal{R}_{13}Z_{\mu})\overline{\chi}\gamma^{\mu}\chi\,$, (2.50)
$\mathcal{L}_{A^{\prime},Z,A_{\gamma}}$ $=\frac{1}{2}\overline{f_{i}}\gamma^{\mu}\left[(v_{i}^{\prime}-a_{i}^{\prime}\gamma^{5})f_{i}A_{\mu}^{\prime}+(v_{i}-a_{i}\gamma^{5})f_{i}Z_{\mu}\right]+eQ_{i}\overline{f_{i}}\gamma^{\mu}f_{i}A_{\gamma\,\mu}\,$, (2.51)
where $eQ_{i}$ is the electric charges of the SM fermions, and
$v_{i}$ $=\left(g_{2}\mathcal{R}_{33}-g_{Y}\mathcal{R}_{23}\right)T_{i}^{3}+2g_{Y}\mathcal{R}_{23}Q_{i}\,$, (2.52)
$a_{i}$ $=\left(g_{2}\mathcal{R}_{33}-g_{Y}\mathcal{R}_{23}\right)T_{i}^{3}\,$, (2.53)
$v_{i}^{\prime}$ $=\left(g_{2}\mathcal{R}_{31}-g_{Y}\mathcal{R}_{21}\right)T_{i}^{3}+2g_{Y}\mathcal{R}_{21}Q_{i}\,$, (2.54)
$a_{i}^{\prime}$ $=\left(g_{2}\mathcal{R}_{31}-g_{Y}\mathcal{R}_{21}\right)T_{i}^{3}\,.$ (2.55)
3 Gauge dependence of the scalar effective action
In gauge theories, the Higgs effective potential as a function of arbitrary background field is not a physical observable and is in general gauge-dependent. Consequently, the effective potential depends not only on the $U(1)_{x}$ parameters $g_{x}$, $\lambda_{x}$, and $v_{x}$, but also on the background field value $\phi_{c}$ and on the gauge-fixing parameter $\xi$. In the conventional approach, one evaluates the gauge-dependent effective potential and the resulting effective action, which leads to gauge-dependent gravitational wave results.
In this section, we review the gauge-independent formulation of the effective action based on the Nielsen identity. This framework is essential for obtaining well-defined gravitational wave predictions from cosmological phase transitions.
3.1 The gauge-dependent scalar effective potential
The full (gauge-dependent) effective potential is given by
$V_{{\rm eff}}(\phi_{c},T,\xi)=V_{0}(\phi_{c})+V_{0}^{{\rm\text{1-loop}}}(\phi_{c},\xi)+V_{{\rm T}}^{{\rm{\rm\text{1-loop}}}}(\phi_{c},T,\xi)+V_{{\rm daisy}}(\phi_{c},T,\xi)\,$, (3.1)
where $V_{0}$ is the tree-level potential, $V_{0}^{{\rm{\rm\text{1-loop}}}}$ and $V_{{\rm T}}^{{\rm{\rm\text{1-loop}}}}$ represent the one-loop corrections at zero and finite temperatures respectively, and $V_{{\rm daisy}}$ corresponds to the daisy contribution arising from higher-loop resummation. The tree-level potential $V_{0}$ is written as
$V_{0}(\phi_{c})=-\frac{\mu_{x}^{2}}{2}\phi_{c}^{2}+\frac{\lambda_{x}}{4}\phi_{c}^{4}\,.$ (3.2)
At zero temperature, the one-loop effective potential after dimensional regularization is expressed as
$V_{0}^{{\rm\text{1-loop}}}(\phi_{c})=\sum_{i}\pm g_{i}\frac{m_{i}^{4}(\phi_{c})}{64\pi^{2}}\Bigl[\log\Bigl(\frac{m_{i}^{2}(\phi_{c})}{\Lambda^{2}}\Bigr)-C_{i}-C_{{\rm UV}}\Bigr]\,$, (3.3)
where the prefactor $\pm 1$ corresponds to bosons and fermions respectively, $g_{i}$ denotes the degree of freedom of particle $i$, and $\Lambda$ is the renormalization scale. In this work, we set the scale $\Lambda$ to be of order the $U(1)_{x}$ gauge boson mass, $\Lambda=g_{x}v_{x}$.
The renormalization constants are $C_{i}=5/6$ for gauge bosons and $C_{i}=3/2$ for scalars and fermions, while $C_{{\rm UV}}=\frac{2}{\epsilon}-\gamma+\log 4\pi$ with $\epsilon=4-D$. It is convenient to adopt the $\overline{\text{MS}}$ renormalization scheme [^3] Alternatively, one may adopt an on-shell renormalization scheme, which is commonly used in supercooled phase transition analyses, by introducing an additional counterterm potential $V_{\rm ct}$, written as $V_{{\rm ct}}(\phi_{c})=-\frac{\delta\mu_{x}^{2}}{2}\phi_{c}^{2}+\frac{\delta\lambda_{x}}{4}\phi_{c}^{4}\,$, (3.4) which cancels the ultraviolet divergence of the one-loop correction at zero temperature. $V_{{\rm ct}}$ is determined by imposing the tadpole and on-shell mass conditions at zero temperature $\frac{\partial(V_{{\rm CW}}+V_{{\rm ct}})}{\partial\phi_{c}}\biggr|_{\phi_{c}=v_{x}}=0\,,\qquad\;\frac{\partial^{2}(V_{{\rm CW}}+V_{{\rm ct}})}{\partial\phi_{c}^{2}}\biggr|_{\phi_{c}=v_{x}}=0\,$, (3.5) ensuring that the potential minimum and the field mass coincide with their tree-level values. However, in this scheme some of the counterterms are $\xi$-dependent and may spoil the power-counting procedure required for the gauge-independent analysis presented in the next subsection. We therefore adopt the $\overline{\rm MS}$ scheme throughout this paper. to remove the ultraviolet divergence, under which the one-loop correction at zero temperature reduces to the Coleman-Weinberg potential, written as
$V_{{\rm CW}}(\phi_{c})=$ $\;\sum_{i}\pm g_{i}\frac{m_{i}^{4}(\phi_{c})}{64\pi^{2}}\Bigl[\log\Bigl(\frac{m_{i}^{2}(\phi_{c})}{\Lambda^{2}}\Bigr)-C_{i}\Bigr]$
$=$ $\;\frac{1}{64\pi^{2}}\biggl[m_{h_{x}}^{4}\Bigl(\log\frac{m_{h_{x}}^{2}}{\Lambda^{2}}-\frac{3}{2}\Bigr)+m_{G}^{4}\Bigl(\log\frac{m_{G}^{2}}{\Lambda^{2}}-\frac{3}{2}\Bigr)$
$+3m_{A^{\prime}}^{4}\Bigl(\log\frac{m_{A^{\prime}}^{2}}{\Lambda^{2}}-\frac{5}{6}\Bigr)-m_{c}^{4}\Bigl(\log\frac{m_{c}^{2}}{\Lambda^{2}}-\frac{3}{2}\Bigr)\biggr]\,.$ (3.6)
In the last line of the above equation, the coefficient 1-1 from the $m_{c}^{4}$ term arises from the +1+1 contribution of the longitudinal (time-like) gauge scalar mode, c.f., Eq. (2.9), together with the 2-2 contribution from the ghost.
The thermal contribution to the one-loop effective potential is given by
$V_{{\rm T}}^{{\rm 1-loop}}(\phi_{c},T)=\sum_{i}\pm\,g_{i}\,\frac{T^{4}}{2\pi^{2}}\,J_{B/F}\biggl(\frac{m_{i}^{2}(\phi_{c})}{T^{2}}\biggr)\,$, (3.7)
where the thermal function $J$ is defined as
$J_{B/F}(x^{2})\equiv\int_{0}^{\infty}{\rm d}y\,y^{2}\log\Bigl[1\mp\exp\bigl(-\sqrt{y^{2}+x^{2}}\bigr)\Bigr]\,.$ (3.8)
In the high-temperature limit ($x=m/T\ll 1$), the thermal function can be expanded as [67]
$J_{B}(x^{2})$ $\approx-\frac{\pi^{4}}{45}+\frac{\pi^{2}}{12}x^{2}-\frac{\pi}{6}x^{3}-\frac{1}{32}x^{4}\log x^{2}+\cdots\,$, (3.9)
$J_{F}(x^{2})$ $\approx\frac{7\pi^{4}}{360}-\frac{\pi^{2}}{24}x^{2}-\frac{1}{32}x^{4}\log x^{2}+\cdots\,.$ (3.10)
In the low-temperature regime ($x\gg 1$), the expansion takes the form [68, 69]
$J_{B/F}(x^{2})\approx\mp\sum_{l=1}^{k}\frac{(\pm 1)^{l}}{l^{2}}x^{2}K_{2}(xl)\,$, (3.11)
where $K_{2}$ is the modified Bessel function of the second kind, and $k$ denotes the order at which the series is truncated. It has been shown that truncating at $k=2,3$ provides a good approximation [68].
Daisy (or ring) resummation [2], which sums the leading infrared-divergent bosonic zero-mode contributions to all orders [^4] Another approach is to integrate out the temperature dependence by matching to a three-dimensional EFT [70, 71, 72, 73, 74, 75]. and thereby renders the finite-temperature effective potential infrared finite in regimes where naive perturbation theory would otherwise break down, must be included in the full thermal Higgs effective potential. Two approaches are commonly used for daisy resummation: the Parwani method [76] and the Arnold-Espinosa method [3]. In the Parwani method, the daisy correction is implemented by replacing the tree-level masses in $J_{B}$ functions with thermal masses,
$m_{i}^{2}(\phi_{c})\to m_{i}^{2}(\phi_{c})+\Pi_{i}(T)\,.$ (3.12)
In the Arnold-Espinosa method that we follow in this work, an additional term is added to the effective potential,
$V_{{\rm daisy}}(\phi_{c},T)=-\frac{T}{12\pi}\sum_{i}g_{i}\Bigl\{\bigl[m_{i}^{2}(\phi_{c})+\Pi_{i}(T)\bigr]^{3/2}-\bigl[m_{i}^{2}(\phi_{c})\bigr]^{3/2}\Bigr\}\,$, (3.13)
where the sum runs over the longitudinal gauge boson and scalar modes. The one-loop thermal masses for the $U(1)_{x}$ dark sector particles are given by [30]
$\Pi_{A^{\prime}}=\frac{1}{3}g_{x}^{2}T^{2}+\sum_{i=1}^{n}\frac{1}{6}g_{x}^{2}Q_{i}^{2}T^{2}\,,\qquad\Pi_{h_{x}}=\frac{1}{3}\lambda_{x}T^{2}+\frac{1}{4}g_{x}^{2}T^{2}\,,\qquad\Pi_{G}=\frac{1}{4}\lambda_{x}T^{2}\,$, (3.14)
where the second term in $\Pi_{A^{\prime}}$ arises from chiral fermion loops. For a Dirac fermion $\chi$ with $U(1)_{x}$ charge $Q_{x}$, it can be treated as two chiral fermions with the same charge, one obtains
$\Pi_{A^{\prime}}=\frac{1}{3}g_{x}^{2}(1+Q_{x}^{2})T^{2}\,.$ (3.15)
The two prescriptions agree at leading infrared order but can differ by higher-order terms. Although the Parwani approach is often numerically smoother, the Arnold–Espinosa prescription is generally preferred for a controlled perturbative expansion and is therefore more suitable for the gauge-independent analysis.
3.2 Gauge-independent effective action
As discussed above, the effective potential evaluated at a fixed background field value is in general gauge-dependent, and therefore cannot by itself be used to define an unambiguous gravitational wave signal. A gauge-independent treatment can be formulated using the Nielsen identity, which states that an infinitesimal change in the gauge parameter $\xi$ can be compensated by a corresponding field redefinition of the background configuration. In this sense, the $\xi$ dependence of $V_{\rm eff}$ at fixed $\phi$ is a coordinate effect in field space. The induced shift of the background field from the field redefinition, is of higher order in the $\hbar$ expansion and therefore does not appear at the leading order.
Gauge independence of physical quantities, such as the bounce action, the nucleation rate, and related observables, follows when they are evaluated on-shell, i.e., on configurations satisfying $\delta S/\delta\phi=0$, within a consistent loop ($\hbar$) expansion. Under this procedure, the relevant effective action is gauge-independent, leading to well-defined gravitational wave predictions for a given model.
In a weakly coupled $U(1)_{x}$ theory, the $\hbar$ order tracks powers of $g_{x}^{2}$ and $\lambda_{x}$. Every gauge interaction brings at least two powers of $g_{x}$ per closed loop in the Abelian $U(1)_{x}$ theory, and scalar self-interactions bring one power of $\lambda_{x}$ at one loop. Hence, each extra loop corresponds to one extra factor of $g_{x}^{2}$ or $\lambda_{x}$. With an appropriate choice of power counting, $\lambda_{x}\sim g_{x}^{\,n}$, which depends on the temperature regime and will be specified in this subsection, one can organize the effective potential as an expansion in powers of the $U(1)_{x}$ gauge coupling $g_{x}$. This ordering separates leading and subleading contributions in a controlled manner and ensures that the resulting effective action is gauge-independent at the order considered, guaranteed by the Nielsen identity.
In this subsection, we review the gauge-independent analysis of the scalar effective action for phase transitions at zero temperature, in the high-temperature regime, and in the low-temperature (supercooled) regime.
3.2.1 The field-dependent effective action
In quantum field theory, the tree-level potential is modified by higher-order radiative corrections, which can be calculated using the functional method. Consider a scalar field $\phi$ coupled to a classical source $J$, the partition function is given by
$Z[J]={\rm e}^{{\rm i}W[J]}=\int\mathcal{D}\phi\,\exp\Bigl\{{\rm i}\bigl[S[\phi]+\int{\rm d}^{4}x\,J\phi\bigr]\Bigr\}\,$, (3.16)
where $S[\phi]=\int{\rm d}^{4}x\,\mathcal{L}[\phi]$ denotes the action. The functional derivative of $W[J]$ with respect to $J(x)$ is computed as
$\frac{\delta W[J]}{\delta J(x)}=-{\rm i}\frac{\delta\log Z[J]}{\delta J(x)}=\frac{\int\mathcal{D}\phi\,{\rm e}^{{\rm i}\int(\mathcal{L}+J\phi)}\phi(x)}{\int\mathcal{D}\phi\,{\rm e}^{{\rm i}\int(\mathcal{L}+J\phi)}}\,$, (3.17)
corresponding to the vacuum expectation value of $\phi(x)$ in the presence of a non-zero external source $J(x)$. This quantity defines the classical field (mean field) $\phi_{c}(x)$ as
$\frac{\delta W[J]}{\delta J(x)}=\bigl\langle\Omega\bigr|\phi(x)\bigr|\Omega\bigr\rangle_{J}\equiv\phi_{c}(x)\,.$ (3.18)
The effective action $\Gamma[\phi]$ is then defined through the Legendre transformation of $W[J]$ as
$\Gamma[\phi_{c}]\equiv W[J]-\int{\rm d}^{4}x\,J(x)\phi_{c}(x)\,$, (3.19)
satisfying
$\frac{\delta}{\delta\phi_{c}(x)}\Gamma[\phi_{c}]=-J(x)\,.$ (3.20)
When the external source $J(x)$ is set to zero, one obtains
$\frac{\delta}{\delta\phi_{c}(x)}\Gamma[\phi_{c}]\Bigr|_{J=0}=0\,.$ (3.21)
For a translation-invariant vacuum state, one may obtain an $x$-independent solution $\phi_{c}$ and express the effective action in form of
$\Gamma[\phi_{c}]=-\int{\rm d}^{4}x\,V_{{\rm eff}}(\phi_{c})\,$, (3.22)
where $V_{{\rm eff}}$ denotes the effective potential. Applying a Fourier transformation from position to momentum space, $\tilde{\phi}(p)=\int{\rm d}^{4}x\,{\rm e}^{-{\rm i}px}\phi(x)$, the effective potential can be recast as
$V_{{\rm eff}}(\phi_{c})=-\sum_{n=0}^{\infty}\,\frac{\phi_{c}^{n}}{n!}\,\Gamma^{(n)}(p=0)\,$, (3.23)
where $\Gamma^{(n)}$ represents the $n$-point effective vertex in momentum space with all external momenta set to zero.
In what follows, we use $\phi_{c}$ to denote the background (mean) field. For a homogeneous configuration, the true vacuum at temperature $T$ is given by the constant field value $\phi_{c}$ that minimizes the finite-temperature effective potential $V_{\rm eff}(\phi_{c},T)$ globally.
3.2.2 Phase transition at zero temperature
The gauge-independent bubble nucleation rate at zero temperature is derived in [37] by employing the power counting of the gauge coupling and a derivative expansion of the Nielsen identity [77]. In this subsection, we review the computation in [37] demonstrating the gauge independence of the effective action at zero temperature, and the extension to finite temperature will be discussed in Section 3.2.3 and 3.2.4.
The gauge dependence of the effective action in $R_{\xi}$ gauges governed by the Nielsen identity is expressed as
$\xi\frac{\partial S_{{\rm eff}}}{\partial\xi}=-\int{\rm d}^{4}x\,\frac{\delta S_{{\rm eff}}}{\delta\phi_{c}(x)}\,C(x)\,$, (3.24)
where $\phi_{c}$ denotes the background field, $C(x)$ is the Nielsen function, given by
$C(x)=\frac{ig_{x}}{2}\int{\rm d}^{4}y\,\bigl\langle\bar{c}(x)G(x)c(y)\times\bigl(\partial_{\mu}A^{\prime\mu}(y)+\sqrt{2}\xi g_{x}\phi_{c}G(y)\bigr)\bigr\rangle\,$, (3.25)
where $g_{x}$ is a gauge coupling, $c,\bar{c}$ and $G$ represent the ghost, anti-ghost and Goldstone boson, respectively. A derivative expansion of $S_{{\rm eff}}$, $C(x)$ and $\delta S_{{\rm eff}}/\delta\phi_{c}$ can then be performed in powers of the derivatives of $\phi_{c}$, yielding
$S_{{\rm eff}}$ $=\int{\rm d}^{4}x\,\bigl[V_{{\rm eff}}(\phi_{c})+\tfrac{1}{2}Z(\phi_{c})(\partial_{\mu}\phi_{c})^{2}+\mathcal{O}(\partial^{4})\bigr]\,$, (3.26)
$C(x)$ $=C_{0}(\phi_{c})+D(\phi)(\partial_{\mu}\phi_{c})^{2}-\partial^{\mu}\bigl[\tilde{D}(\phi_{c})\partial_{\mu}\phi_{c}\bigr]+\mathcal{O}(\partial^{4})\,$, (3.27)
$\frac{\delta S_{{\rm eff}}}{\delta\phi_{c}}$ $=\frac{\partial V_{{\rm eff}}(\phi_{c})}{\partial\phi_{c}}+\frac{1}{2}\frac{\partial Z}{\partial\phi_{c}}(\partial_{\mu}\phi_{c})^{2}-\partial^{\mu}\bigl[Z(\phi_{c})\partial_{\mu}\phi_{c}\bigr]+\mathcal{O}(\partial^{4})\,$, (3.28)
where $Z(\phi_{c})$ denotes the field renormalization factor and $C_{0},D,\tilde{D}$ are Nielsen coefficients; the total derivative term containing $\tilde{D}$ are discussed in [39]. Substituting the derivative expansions given in Eqs. (3.26)–(3.28) into Eq. (3.24), the Nielsen identity at order $\partial^{0}$ and $\partial^{2}$ is derived as
$\xi\frac{\partial V_{{\rm eff}}}{\partial\xi}$ $=-C_{0}\frac{\partial V_{{\rm eff}}}{\partial\phi_{c}}\,$, (3.29)
$\xi\frac{\partial Z}{\partial\xi}$ $=-C_{0}\frac{\partial Z}{\partial\phi_{c}}-2Z\frac{\partial C_{0}}{\partial\phi_{c}}+2D\frac{\partial V_{{\rm eff}}}{\partial\phi_{c}}+2\tilde{D}\frac{\partial^{2}V_{{\rm eff}}}{\partial\phi_{c}^{2}}\,$, (3.30)
where integration by parts is applied and the surface term vanishes as follows:
$\int{\rm d}^{4}x\,C_{0}(\phi_{c})\partial_{\mu}\bigl[Z(\phi_{c})\partial_{\mu}\phi_{c}\bigr]$ $=\int{\rm d}^{4}x\,\Bigl\{\partial_{\mu}\bigl[C_{0}(\phi_{c})Z(\phi_{c})\partial_{\mu}\phi_{c}\bigr]-\bigl[\partial_{\mu}C_{0}(\phi_{c})\bigr]Z(\phi_{c})\partial_{\mu}\phi_{c}\Bigr\}$
$=\int{\rm d}^{4}x\,\bigl[\partial_{\mu}C_{0}(\phi_{c})\bigr]Z(\phi_{c})\partial_{\mu}\phi_{c}$
$=\int{\rm d}^{4}x\,Z(\phi_{c})\,\frac{\partial C_{0}}{\partial\phi_{c}}\,(\partial_{\mu}\phi_{c})^{2}\,.$ (3.31)
Now take the loop ($\hbar$) expansion to the potential and the field strength renormalization as
$V=V_{0}+\hbar V_{1}+\hbar^{2}V_{2}+\cdots\,,\qquad Z=1+\hbar Z_{1}+\hbar^{2}Z_{2}+\cdots\,.$ (3.32)
In a weakly coupled $U(1)_{x}$ theory, the order in $\hbar$ tracks powers of $g_{x}^{2}$ and $\lambda_{x}$. The analysis below will proceed by expanding $V$ and $Z$, as well as the Nielsen coefficients, in powers of the $U(1)_{x}$ gauge coupling $g_{x}$.
For the zero-temperature analysis, a Higgs-like scalar potential without thermal effects does not generically produce a phase transition. A first-order transition requires two minima separated by a barrier, with the true vacuum lying below the false vacuum initially located at $\phi=0$. One simple way to realize this in a toy setup is to choose a positive mass term for the scalar,
$V(\phi)\supset\frac{1}{2}\,\mu_{x}^{2}\phi^{2}\,,\qquad\mu_{x}^{2}>0$ (3.33)
in contrast to the usual Higgs potential with a negative mass-squared term. Additionally, since the barrier is generated by the Coleman-Weinberg contribution, the positive mass term $\sim\mu^{2}$ cannot be parametrically large, otherwise the potential will not develop. This setup can thus support a zero-temperature phase transition and is often adopted for pioneering studies [37].
For a gauge theory in which the gauge symmetry is broken by a Higgs field, a crucial prerequisite for a controlled expansion is to choose an appropriate power counting for the Higgs quartic coupling $\lambda_{x}$ in terms of the gauge coupling $g_{x}$. In the zero-temperature regime, one commonly adopts the scaling
$\lambda_{x}\sim g_{x}^{4}\,.$ (3.34)
This scaling has two key insights:
  • The generation of the barrier at zero temperature: In a gauged $U(1)_{x}$ theory, the dominant Coleman-Weinberg contribution scales as
    $V_{\rm CW}(\phi)\sim\frac{g_{x}^{4}}{16\pi^{2}}\phi^{4}\log\frac{g_{x}^{2}\phi^{2}}{\Lambda^{2}}\sim\frac{g_{x}^{4}}{16\pi^{2}}\phi^{4}\,.$ (3.35)
    The extremum condition $V^{\prime}(v_{x})=0$ implies
    $\mu_{x}^{2}\sim v_{x}^{2}\left(\lambda_{x}+\frac{g_{x}^{4}}{16\pi^{2}}\right)\,.$ (3.36)
    For the Coleman-Weinberg contribution to compete with the tree-level quartic term and thereby generate a barrier, one requires $\lambda_{x}\sim g_{x}^{4}$. It then follows that $\mu_{x}^{2}\sim g_{x}^{4}\phi_{c}^{2}$.
  • The validity of the derivative expansion: The derivative expansion is a low-momentum expansion obtained after integrating out the heavy modes of the theory, and it is expanded in powers of the ratio of the characteristic momentum scale associated with spatial variations of the background field, $k_{\rm field}$, to the mass scale of heavy modes that have been integrated out, $\Lambda(\phi_{\rm wall})$. Here $\phi_{\rm wall}$ denotes the typical field value within the bubble wall. The derivative expansion is thus valid when $k_{\rm field}/\Lambda(\phi_{\rm wall})\ll 1$, i.e., $k_{\rm field}\ll\Lambda(\phi_{\rm wall})$.
    The characteristic momentum is estimated as
    $k_{\rm field}\sim\frac{1}{L_{\rm wall}}\sim\sqrt{V^{\prime\prime}(\phi_{\rm wall})}\,$, (3.37)
    where $L_{\rm wall}$ is the characteristic thickness of the bubble wall. For the zero-temperature phase transition,
    $k_{\rm field}\sim\frac{g_{x}^{2}}{4\pi}\,\phi_{\rm wall}\,$, (3.38)
    which is parametrically much smaller than the scale at which the dark photon has been integrated out, $\Lambda=g_{x}\,\phi_{\rm wall}$, and thus the derivative expansion is valid at zero temperature.
At zero temperature, the effective potential $V_{{\rm eff}}$ is given by
$V^{0{\rm T}}_{{\rm eff}}=V^{0{\rm T}}_{g_{x}^{4}}+V^{0{\rm T}}_{g_{x}^{6}}+\mathcal{O}(g_{x}^{8})\,$, (3.39)
with the leading and subleading effective potentials written as
$V^{0{\rm T}}_{g_{x}^{4}}$ $=\frac{\mu_{x}^{2}}{2}\phi_{c}^{2}+\frac{\lambda_{x}}{4}\phi_{c}^{4}+3\,\frac{m_{A^{\prime}}^{4}}{64\pi^{2}}\Bigl(\log\frac{m_{A^{\prime}}^{2}}{\Lambda^{2}}-\frac{5}{6}\Bigr)\,$, (3.40)
$V^{0{\rm T}}_{g_{x}^{6}}$ $=\frac{m_{G}^{4}}{64\pi^{2}}\Bigl(\log\frac{m_{G}^{2}}{\Lambda^{2}}-\frac{3}{2}\Bigr)-\frac{m_{c}^{4}}{64\pi^{2}}\Bigl(\log\frac{m_{c}^{2}}{\Lambda^{2}}-\frac{3}{2}\Bigr)\,$, (3.41)
The field strength renormalization $Z$ is expanded in powers of the gauge coupling $g_{x}$ as
$Z$ $=1+Z_{g_{x}^{2}}+\mathcal{O}(g_{x}^{4})\,.$
$=\frac{g_{x}^{2}}{16\pi^{2}}\Bigl(\xi\log\frac{m_{c}^{2}}{\Lambda^{2}}+3\log\frac{m_{A^{\prime}}^{2}}{\Lambda^{2}}+\xi\Bigr)\,$, (3.42)
and perturbation analysis shows that the Nielson coefficient $C_{0}(\phi_{c})$ starts at order $g_{x}^{2}$ and $D,\tilde{D}$ starts at unity, expressed as
$C_{0}$ $=C_{g_{x}^{2}}+\mathcal{O}(g_{x}^{4})$
$=-\frac{\xi g_{x}^{2}\phi_{c}}{32\pi^{2}}\,\log\frac{m_{c}^{2}}{\Lambda^{2}}+\mathcal{O}(g_{x}^{4})\,$, (3.43)
$D,\tilde{D}$ $=1+\mathcal{O}(g_{x}^{2})\,.$ (3.44)
The leading and subleading contributions to the Nielsen identity are thus obtained as follows
$\xi\frac{\partial V^{0{\rm T}}_{g_{x}^{4}}}{\partial\xi}$ $=0\,$, (3.45)
$\xi\frac{\partial V^{0{\rm T}}_{g_{x}^{6}}}{\partial\xi}$ $=-C_{g_{x}^{2}}\frac{\partial V^{0{\rm T}}_{g_{x}^{4}}}{\partial\phi_{c}}\,$, (3.46)
$\xi\frac{\partial Z_{g_{x}^{2}}}{\partial\xi}$ $=-2\frac{\partial C_{g_{x}^{2}}}{\partial\phi_{c}}\,$, (3.47)
which hold analytically.
The standard procedure for calculating the bubble nucleation rate at zero temperature is finding a bounce solution $\phi_{b}$ [78, 79] from the leading-order equation of motion
$\Box\phi_{b}=\frac{\partial V^{0{\rm T}}_{g_{x}^{4}}}{\partial\phi_{c}}\biggr|_{\phi_{c}=\phi_{b}}\,.$ (3.48)
The nucleation rate is then obtained as
$\Gamma=\mathcal{A}\,{\rm e}^{-(S^{0{\rm T}}_{0}+S^{0{\rm T}}_{1})}\,$, (3.49)
where the coefficient $\mathcal{A}$ is expressed in terms of functional determinants associated with the characteristic mass scales of the theory, $S^{0{\rm T}}_{0}$ and $S^{0{\rm T}}_{1}$ denote the Euclidean action of the bounce solution, given by
$S^{0{\rm T}}_{0}$ $=\int{\rm d}^{4}x\,\Bigl[V^{0{\rm T}}_{g_{x}^{4}}(\phi_{b})+\frac{1}{2}(\partial_{\mu}\phi_{b})^{2}\Bigr]\,$, (3.50)
$S^{0{\rm T}}_{1}$ $=\int{\rm d}^{4}x\,\Bigl[V^{0{\rm T}}_{g_{x}^{6}}(\phi_{b},\xi)+\frac{1}{2}Z_{g_{x}^{2}}(\phi_{b},\xi)\,(\partial_{\mu}\phi_{b})^{2}\Bigr]\,.$ (3.51)
$S_{0}$ is gauge-independent since $V^{0{\rm T}}_{g_{x}^{4}}$ is $\xi$-independent. Using Eqs. (3.41) and (3.42), one finds
$\xi\frac{\partial S^{0{\rm T}}_{1}}{\partial\xi}$ $=\int{\rm d}^{4}x\,\Bigl[\xi\frac{\partial V^{0{\rm T}}_{g_{x}^{6}}}{\partial\xi}+\frac{1}{2}\xi\frac{\partial Z_{g_{x}^{2}}}{\partial\xi}\,(\partial_{\mu}\phi_{b})^{2}\Bigr]$
$=-\int{\rm d}^{4}x\,\Bigl[C_{g_{x}^{2}}\frac{\partial V^{0{\rm T}}_{g_{x}^{4}}}{\partial\phi_{c}}+\frac{\partial C_{g_{x}^{2}}}{\partial\phi_{c}}\,(\partial_{\mu}\phi_{b})^{2}\Bigr]\biggr|_{\phi_{c}=\phi_{b}}$
$=-\int{\rm d}^{4}x\Bigl[C_{g_{x}^{2}}\frac{\partial V^{0{\rm T}}_{g_{x}^{4}}}{\partial\phi_{c}}+(\partial_{\mu}C_{g_{x}^{2}})(\partial_{\mu}\phi_{b})\Bigr]\biggr|_{\phi_{c}=\phi_{b}}$
$=-\int{\rm d}^{4}x\,C_{g_{x}^{2}}\Bigl[\frac{\partial V^{0{\rm T}}_{g_{x}^{4}}}{\partial\phi_{c}}-\Box\phi_{b}\Bigr]\biggr|_{\phi_{c}=\phi_{b}}$
$=0\,$, (3.52)
the bounce solution $\phi_{b}$ ensures the gauge independence of the subleading effective action $S^{0{\rm T}}_{1}$.
3.2.3 Phase transition at high temperatures ($T\gg m_{A^{\prime}}$)
High-temperature phase transitions typically occur at temperatures only slightly below the critical temperature $T_{c}$, which is generally much larger than the $U(1)_{x}$ gauge boson mass. In this situation, the percolation temperature that characterizes the onset of the phase transition is usually close to the nucleation temperature, $T_{p}\approx T_{n}$, as will be discussed in Section 4.2. At $T_{n}$, the background field value satisfies $\phi_{c}\gtrsim v_{x}$, and the transition temperature is much larger than the particle masses in the $U(1)_{x}$ sector, i.e. $T_{n}\gg m_{A^{\prime}}$. In this regime, the calculation can be simplified by expanding in $m/T$ and neglecting higher-order terms.
The bubble nucleation rate at finite temperature is given by
$\Gamma(T)$ $=\mathcal{A}(T)\,{\rm e}^{-S_{3}(T)/T}\,$, (3.53)
$S_{3}(T)$ $=\int{\rm d}^{3}x\,\bigl[V_{{\rm eff}}(\phi_{c},T)+\tfrac{1}{2}Z(\phi_{c},T)(\partial_{i}\phi_{c})^{2}+\mathcal{O}(\partial^{4})\bigr]\,$, (3.54)
where $\phi_{c}$ denotes the background field, $Z(\phi_{c},T)$ is the field renormalization factor, and the prefactor $\mathcal{A}(T)$ will be given in Eq. (4.4).
We now implement the power counting procedure for the high-temperature phase transition, following [39, 43, 44]. The nucleation temperature $T_{n}$ marks the onset of the phase transition and lies slightly below the critical temperature $T_{c}$, at which the two minima of the effective potential become degenerate. Therefore $T_{c}$ can be taken as a representative characteristic temperature of the transition. Including the one-loop thermal corrections to $V_{{\rm eff}}$ and employing the high-temperature expansion in Eq. (3.9), the leading contribution to the effective potential becomes
$V_{{\rm LO}}=$ $-\frac{\mu_{x}^{2}}{2}\phi_{c}^{2}+\frac{\lambda_{x}}{4}\phi_{c}^{4}+\frac{T^{2}}{24}(m_{h_{x}}^{2}+3m_{A^{\prime}}^{2}+m_{G}^{2}-m_{c}^{2})+\cdots$
$=$ $-\frac{1}{2}\mu_{{\rm eff}}^{2}\phi_{c}^{2}+\frac{\lambda_{x}}{4}\phi_{c}^{4}\cdots\,$, (3.55)
where particle masses are given in Eqs. (2.9)–(2.12), and
$\mu_{{\rm eff}}^{2}=-\mu_{x}^{2}+\bigl(\tfrac{1}{3}\lambda_{x}+\tfrac{1}{4}g_{x}^{2}\bigr)T^{2}\,.$ (3.56)
The negative mass term $\mu_{x}^{2}$ competes with the positive $T^{2}$ corrections, leading to a critical crossover temperature $T_{o}$ where the Higgs effective mass is zero, i.e., $\mu_{{\rm eff}}^{2}=0$. This temperatures is just slightly below $T_{c}$ and a little bit above $T_{n}$, i.e., $T_{c}\gtrsim T_{o}\gtrsim T_{n}$. Near the high-temperature phase transition, all terms in the potential are expected to be of the same order of magnitude [3, 42], [^5] The full thermal potential can be effectively written as $V(\phi,T)=D(T^{2}-T_{0}^{2})\phi^{2}-ET\phi^{3}+\frac{\lambda_{x}}{4}\phi^{4}\,$, (3.57) where $T_{0}$ is the temperature at which $\phi=0$ is no longer a local minimum. At the critical temperature $T_{c}$, $V^{\prime}(\phi,T_{c})=0\,,\qquad V(\phi,T_{c})=V(0,T_{c})\,$, (3.58) which solve to $\phi_{c}=\frac{2E}{\lambda_{x}}T_{c}\,,\qquad D(T^{2}-T_{0}^{2})=\frac{E^{2}}{\lambda_{x}}T_{c}^{2}\,,\qquad{\rm at}\ T_{c}\,.$ (3.59) Plugging in back to the potential Eq. (3.57) we find all three terms are the same parametric size $\sim E^{4}T_{c}^{4}/\lambda_{x}^{3}$. At $T_{n}$ which lies just slightly below $T_{c}$, relation Eq. (3.59) approximately holds and thus $\frac{\phi_{c}}{T_{n}}\approx\frac{2E}{\lambda_{x}}\sim\mathcal{O}(1)\,.$ (3.60) In gauge theory, $E\sim g_{x}^{3}$ and thus $\lambda_{x}\sim g_{x}^{3}$. yielding the following power counting relation:
$\lambda_{x}\sim g_{x}^{3}\,,\qquad\mu_{{\rm eff}}^{2}\sim g_{x}^{3}T^{2}\,,\qquad T\sim\phi_{c}\,.$ (3.61)
For the scaling adopted above, the derivative expansion expanded in powers of $k_{\rm field}/\Lambda$, where the characteristic momentum scale associated with field variations is $k_{\rm field}\sim\sqrt{V^{\prime\prime}}\sim g^{3/2}\phi_{c}$, and the cutoff scale $\Lambda$ is set by the non-zero Matsubara modes, $\Lambda\sim 2\pi T$. Hence,
$\frac{k_{\rm field}}{\Lambda}\sim\frac{g_{x}^{3/2}}{2\pi}\ll 1\,$, (3.62)
ensuring that the derivative expansion is well controlled in the high-temperature regime.
In practice, our parameter scan shows that a high-temperature phase transition can also occur for $\lambda_{x}\ll g_{x}^{3}$. In this case, the nucleation temperature $T_{n}$ can lie well below $T_{c}$, while the high-temperature expansion with $m_{A^{\prime}}/T_{n}\ll 1$ remains reasonably applicable.
With the thermal corrections, the corrected field-dependent masses become
$m_{A^{\prime}}^{2}$ $=g_{x}^{2}\phi_{c}^{2}\,,\quad m_{A^{\prime}_{0}}^{2}=\xi g_{x}^{2}\phi_{c}^{2}\,,\quad m_{c}^{2}=\xi g_{x}^{2}\phi_{c}^{2}\,$, (3.63)
$m_{h_{x}}^{2}$ $=-\mu_{{\rm eff}}^{2}+3\lambda_{x}\phi_{c}^{2}\,$, (3.64)
$m_{G}^{2}$ $=-\mu_{{\rm eff}}^{2}+\lambda_{x}\phi_{c}^{2}+\xi g_{x}^{2}\phi_{c}^{2}\,.$ (3.65)
Using the power counting introduced in Eq. (3.61), the leading-order effective potential is obtained as
$V^{\rm HT}_{g_{x}^{3}}=$ $-\frac{\mu_{x}^{2}}{2}\phi_{c}^{2}+\frac{\lambda_{x}}{4}\phi_{c}^{4}+\frac{T^{2}}{24}(m_{h_{x}}^{2}+3m_{A^{\prime}}^{2}+m_{G}^{2}-m_{c}^{2})$
$-\frac{T}{12}\bigl[2m_{A^{\prime}}^{3}+(m_{A^{\prime}}^{2}+\Pi_{A^{\prime}})^{3/2}\bigr]$
$=$ $-\frac{1}{2}\mu_{{\rm eff}}^{2}\phi_{c}^{2}+\frac{\lambda_{x}}{4}\phi_{c}^{4}-\frac{g_{x}^{3}T}{12\pi}\bigl[2\phi_{c}^{3}+\bigl(\tfrac{1}{3}T^{2}+\phi_{c}^{2}\bigr)^{3/2}\bigr]\,.$ (3.66)
To evaluate the next-to-leading order contribution, we first examine the Coleman-Weinberg potential, c.f., Eq. (3.6),
$V_{{\rm CW}}=F(m_{h_{x}}^{2})+3F(m_{A^{\prime}}^{2})+F(m_{G}^{2})-F(m_{c}^{2})\,$, (3.67)
where
$F(m_{i}^{2})\equiv\frac{m_{i}^{4}}{64\pi^{2}}\Bigl(\log\frac{m_{i}^{2}}{\Lambda^{2}}-C_{i}\Bigr)\,.$ (3.68)
Among these terms, the gauge-independent contributions are $F(m_{h_{x}}^{2})\sim\mathcal{O}(g_{x}^{6})$ and $F(m_{A^{\prime}}^{2})\sim\mathcal{O}(g_{x}^{4})$. We then perform a Taylor expansion $f(x)=f(a)+f^{\prime}(a)\,(x-a)+\cdots$ around $m_{c}^{2}$, and we obtain
$F(m_{G}^{2})-F(m_{c}^{2})\simeq F^{\prime}(m_{c}^{2})\,(m_{G}^{2}-m_{c}^{2})\sim\xi\,\mathcal{O}(g_{x}^{5})\,$, (3.69)
where $F^{\prime}(m_{c}^{2})\sim\mathcal{O}(m_{c}^{2})$. The remaining terms of the effective potential are
$\frac{T}{12\pi}\bigl[(m_{h_{x}}^{2}+\Pi_{h_{x}})^{3/2}+(m_{G}^{2}+\Pi_{G})^{3/2}-m_{G}^{3}+\underline{m_{G}^{3}-m_{c}^{3}}\bigr]\,$, (3.70)
where the first two $3/2$-power terms arise from daisy resummation, while the two underlined terms originate from the high-temperature expansion of the bosonic thermal function $J_{B}$, c.f., Eq. (3.9). The term $(m_{h_{x}}^{2}+\Pi_{h_{x}})^{3/2}\sim\mathcal{O}(g_{x}^{9/2})$ is of higher order in the power counting than $\mathcal{O}(g_{x}^{4})$ and is therefore neglected. The gauge-dependent part is evaluated as
$(m_{G}^{2}+\Pi_{G})^{3/2}-m_{c}^{3}$ $=\left(-\mu_{{\rm eff}}^{2}+\lambda_{x}\phi_{c}^{2}+\xi g_{x}^{2}\phi_{c}^{2}+\frac{1}{4}\lambda_{x}T^{2}\right)^{3/2}-\left(\xi g_{x}^{2}\phi_{c}^{2}\right)^{3/2}$
$\approx\left(-\mu_{{\rm eff}}^{2}+\lambda_{x}\phi_{c}^{2}+\xi g_{x}^{2}\phi_{c}^{2}\right)^{3/2}-\left(\xi g_{x}^{2}\phi_{c}^{2}\right)^{3/2}\sim\sqrt{\xi}\,\mathcal{O}(g_{x}^{4})\,$, (3.71)
where $m_{c}^{2}=\xi g_{x}^{2}\phi_{c}^{2}$ is much larger than $\mu_{{\rm eff}}^{2}+\lambda_{x}\phi_{c}^{2}+\frac{1}{4}\lambda_{x}T^{2}$, for $\xi\gtrsim g_{x}$. For the limit $\xi\ll g_{x}$, the gauge-dependent contribution becomes negligible. In either case, this contribution is counted as a second-order effect and is therefore included in the $\mathcal{O}(g_{x}^{4})$ potential, $V_{g_{x}^{4}}$.
Therefore, the next-to-leading order potential is written as
$V^{0}_{g_{x}^{4}}=3\frac{m_{A^{\prime}}^{4}}{64\pi^{2}}\Bigl(\log\frac{m_{A^{\prime}}^{2}}{\Lambda^{2}}-\frac{5}{6}\Bigr)-\frac{T}{12\pi}\left(m_{G}^{3}-m_{c}^{3}\right)\,.$ (3.72)
It was pointed out in [37] that, under a specific choice of power counting, certain higher-loop diagrams can yield sizable contributions. In particular, diagrams containing transverse photon loops along a scalar propagator do not increase the overall parametric order of the amplitude, and therefore may contribute as the same order as lower-loop terms. These contributions can be summed by replacing the tree-level Goldstone mass with a dressed mass [37],
$m_{G}^{2}\to\widetilde{m}_{G}^{2}\equiv\frac{1}{\phi_{c}}\frac{\partial V^{\rm HT}_{g_{x}^{3}}}{\partial\phi_{c}}+\xi g_{x}^{2}\phi_{c}^{2}\,.$ (3.73)
After the substitution, the subleading potential is written as
$\widetilde{V}_{g_{x}^{4}}^{0}=$ $\ 3\frac{m_{A^{\prime}}^{4}}{64\pi^{2}}\Bigl(\log\frac{m_{A^{\prime}}^{2}}{\Lambda^{2}}-\frac{5}{6}\Bigr)-\frac{T}{12\pi}\left(\widetilde{m}_{G}^{3}-m_{c}^{3}\right)$
$\simeq$ $\ 3\frac{m_{A^{\prime}}^{4}}{64\pi^{2}}\Bigl(\log\frac{m_{A^{\prime}}^{2}}{\Lambda^{2}}-\frac{5}{6}\Bigr)-\frac{Tg_{x}\phi_{c}\sqrt{\xi}}{8\pi}\Bigl[-\mu_{{\rm eff}}^{2}+\lambda_{x}\phi_{c}^{2}-\frac{g_{x}^{3}T}{4\pi}\Bigl(2\phi_{c}+\sqrt{\tfrac{1}{3}T^{2}+\phi_{c}^{2}}\Bigr)\Bigr]\,.$ (3.74)
For nucleation-related observables, one typically requires higher accuracy than that provided by the leading-order barrier alone. The next-to-leading order (NLO) effective potential is derived within a three-dimensional finite-temperature effective field theory (EFT) framework. The three-dimensional effective potential is computed in a constant background field $\phi_{3}$ using three-dimensional perturbation theory: the leading-order piece is the tree-level potential plus the leading three-dimensional one-loop contributions, while the NLO piece consists of the relevant two-loop vacuum diagrams in the three-dimensional EFT together with the corrections induced by NLO-matched parameters. After renormalization and consistent power counting, one arrives at [43, 44]
$V^{\rm HT}_{g_{x}^{4}}$ $=-\frac{Tg_{x}\phi_{c}\sqrt{\xi}}{8\pi}\Bigl[-\mu_{{\rm eff}}^{2}+\lambda_{x}\phi_{c}^{2}-\frac{g_{x}^{3}T}{4\pi}\Bigl(2\phi_{c}+\sqrt{\tfrac{1}{3}T^{2}+\phi_{c}^{2}}\Bigr)\Bigr]$
$+\frac{1}{(4\pi)^{2}}g_{x}^{4}T^{2}\phi_{c}^{2}\Bigl[-\frac{3}{2}+\ln\Bigl(\frac{4g_{x}^{2}\phi_{c}^{2}}{\Lambda^{2}}\Bigr)+\frac{1}{2}\ln\Bigl(\frac{4g_{x}^{2}(\tfrac{1}{3}T^{2}+\phi_{c}^{2})}{\Lambda^{2}}\Bigr)\Bigr]\,$, (3.75)
which will be used for the high-temperature phase transition analysis in the following sections.
With the above discussion, the gauge-independent effective action can be thus constructed. The effective potential and the field renormalization factor are expanded in powers of the coupling $g_{x}$ as
$V^{\rm HT}_{{\rm eff}}$ $=V^{\rm HT}_{g_{x}^{3}}+V^{\rm HT}_{g_{x}^{4}}+\mathcal{O}(g_{x}^{9/2}T^{4})\,$, (3.76)
$Z$ $=1+Z_{g_{x}}+\mathcal{O}(g_{x}^{3/2})\>,$ (3.77)
where $Z_{g_{x}}$ is given in [39] as
$Z_{g_{x}}=\frac{g_{x}T}{48\pi}\biggl[-\frac{22}{\phi_{c}}+\frac{\phi_{c}^{2}}{\bigl(\frac{1}{3}T^{2}+\phi_{c}^{2}\bigr)^{3/2}}\biggr]\,.$ (3.78)
The leading-order potential $V_{g_{x}^{3}}$, c.f., Eq. (3.66), is manifestly gauge-independent, and the corresponding bounce solution $\phi_{b}$ satisfies
$\nabla^{2}\phi_{b}=\frac{\partial V_{g_{x}^{3}}}{\partial\phi_{c}}\biggr|_{\phi_{c}=\phi_{b}}\,,\qquad\phi_{b}(\infty)=0\,,\qquad\phi_{b}^{\prime}(0)=0\,.$ (3.79)
The nucleation rate then takes the form
$\Gamma$ $=\mathcal{A}\,{\rm e}^{-(S^{\rm HT}_{0}+S^{\rm HT}_{1})/T}\,$, (3.80)
$S^{\rm HT}_{0}$ $=\int{\rm d}^{3}x\,\Bigl[V^{\rm HT}_{g_{x}^{3}}(\phi_{b})+\frac{1}{2}(\partial_{i}\phi_{b})^{2}\Bigr]\,$, (3.81)
$S^{\rm HT}_{1}$ $=\int{\rm d}^{3}x\,\Bigl[V^{\rm HT}_{g_{x}^{4}}(\phi_{b})+\frac{1}{2}Z_{g_{x}}(\partial_{i}\phi_{b})^{2}\Bigr]\,.$ (3.82)
The leading order effective action is gauge-independent, since the leading order effective potential $V_{g_{x}^{3}}$ is constructed to be $\xi$-independent.
To analyze the gauge dependence of $V_{g_{x}^{4}}$ and $S_{1}$, we employ the finite-temperature Nielsen identity and expand its coefficients
$C=C_{g_{x}}+\mathcal{O}(g_{x}^{3/2})\,,\quad D,\tilde{D}=\mathcal{O}(g_{x}^{-1})\,$, (3.83)
which yields
$\xi\frac{\partial V^{\rm HT}_{g_{x}^{4}}}{\partial\xi}$ $=-C_{g_{x}}\frac{\partial V^{\rm HT}_{g_{x}^{3}}}{\partial\phi_{c}}\,$, (3.84)
$\xi\frac{\partial Z_{g_{x}}}{\partial\xi}$ $=-2\frac{\partial C_{g_{x}}}{\partial\phi_{c}}\,$, (3.85)
where the leading coefficient $C_{g_{x}}=\sqrt{\xi}g_{x}T/(16\pi)$ [39] is independent of the background field $\phi_{c}$, i.e., $\partial C_{g_{x}}/\partial\phi_{c}=0$. Together with the gauge-independent renormalization factor $Z_{g_{x}}$ in Eq. (3.78), such that $\partial Z_{g_{x}}/\partial\xi=0$, they ensure consistency with the Nielsen relation in Eq. (3.85). Finally, applying Gauss’s theorem and the asymptotic behavior of $\phi_{b}$, one arrives
$\xi\frac{\partial S^{\rm HT}_{1}}{\partial\xi}$ $=\int{\rm d}^{3}x\,\Bigl[\xi\frac{\partial V^{\rm HT}_{g_{x}^{4}}}{\partial\xi}+\frac{1}{2}\xi\frac{\partial Z_{g_{x}}}{\partial\xi}(\partial_{i}\phi_{b})^{2}\Bigr]$
$=\int{\rm d}^{3}x\,\Bigl[-C_{g_{x}}\frac{\partial V^{\rm HT}_{g_{x}^{3}}}{\partial\phi_{c}}+0\Bigr]\biggr|_{\phi_{c}=\phi_{b}}$
$=-C_{g_{x}}\int{\rm d}^{3}x\,\nabla^{2}\phi_{b}$
$=-C_{g_{x}}\int{\rm d}^{2}S\,\partial\phi_{b}$
$=0\,.$ (3.86)
In the high-temperature regime, Eqs. (3.84) and (3.85) are satisfied explicitly by the analytic expressions given above, demonstrating that the resulting high-temperature effective action is analytically gauge-independent.
3.2.4 Phase transition at low temperatures ($T\ll m_{A^{\prime}}$)
The supercooled phase transition typically takes place at $T_{p}\ll m_{A^{\prime}}$. In this regime, the nontrivial temperature-dependent contributions relevant to the transition are generally subdominant to the pure field-dependent terms, since $T_{p}\ll\phi_{c}(T_{p})\sim v_{x}$. In this case, however, unlike in the high-temperature regime, the low-temperature thermal contribution does not admit a well-defined analytic expansion in the small parameter $T/m$. Moreover, the thermal effects are significant only when the field value approaches zero. The basic strategy is therefore to treat the thermal corrections as small perturbations added to the zero-temperature effective potential, as was done in [34].
Unlike the zero-temperature toy setup with a positive Higgs mass term, a minimal gauged $U(1)$ Higgs sector contains the standard tachyonic mass-squared term. In the supercooled regime, thermal effects, although being small since $T_{p}\ll v_{x}$, can lift the effective mass term positive and thereby create and maintain a false vacuum at $\phi=0$, as long as the tachyonic mass-squared parameter $\mu_{x}^{2}$ is parametrically small.
In addition to turning the effective mass term positive, a successful supercooled phase transition requires thermal effects to generate a local barrier near $\phi=0$. This barrier arises from the one-loop thermal contribution together with daisy resummation, through the thermal cubic term $\sim T\,\phi^{3}$. Although this term is parametrically suppressed in the low-temperature regime since $T_{p}\ll v_{x}$, it can still induce a barrier in the vicinity of $\phi=0$, where $T_{p}\sim g_{x}\phi$. Consequently, the thermally generated barrier peaks at field values of order $\phi_{\rm barrier}\sim{T_{p}}/{g_{x}}$ in the supercooled phase transition.
The thermal cubic term $\sim T\,\phi^{3}$ responsible for generating the barrier is loop induced and is parametrically suppressed, especially in the low-temperature regime. For this term to compete with the quartic contribution, the quartic coupling $\lambda_{x}$ must itself be loop sized. In particular, $\lambda_{x}$ should not exceed the size implied by the Coleman-Weinberg contribution:
$\frac{\lambda_{x}}{4}\phi^{4}\lesssim\frac{m_{A^{\prime}}^{4}}{64\pi^{2}}\log\frac{m_{A^{\prime}}^{2}}{\Lambda^{2}}\,$, (3.87)
which implies
$\lambda_{x}\lesssim\mathcal{O}\left(\frac{g_{x}^{4}}{16\pi^{2}}\right)\,.$ (3.88)
For this scaling, the characteristic momentum scale associated with spatial variations of the background field can be estimated as
$k_{\rm field}\sim\sqrt{\lambda_{x}}\,\phi_{\rm wall}\sim\frac{g_{x}^{2}}{4\pi}\,\phi_{\rm wall}\,$, (3.89)
whereas the heavy scale that has been integrated out is set by the mass of the $U(1)_{x}$ gauge boson, $\Lambda\sim g_{x}\phi_{\rm wall}$. This implies
$\frac{k_{\rm field}}{\Lambda}\sim\frac{g_{x}}{4\pi}\ll 1\,$, (3.90)
and thus the derivative expansion remains parametrically controlled in the low-temperature regime, even in the $\phi\to 0$ region.
Thus for the supercooled phase transition, we use the scaling [34] to organize the field-dependent terms in the effective potential
$\lambda_{x}\sim g_{x}^{4}\,,\qquad\mu_{x}^{2}\sim g_{x}^{4}\phi_{c}^{2}\,$, (3.91)
For the thermal contributions, as $T\to 0$, the daisy contribution scales as $V_{\rm daisy}\sim T\,\phi_{c}^{3}$, while the one-loop thermal contribution is further suppressed exponentially,
$\frac{T^{4}}{2\pi^{2}}\,J_{B}\!\left(\frac{m^{2}}{T^{2}}\right)\sim T^{4}\,{\rm e}^{-m/T}\,.$ (3.92)
Therefore, away from the barrier region, thermal effects are generally negligible because $T\ll\phi_{c}(T)$, and all thermal contributions can be safely neglected relative to the purely field-dependent terms.
Nevertheless, thermal effects are essential for generating the barrier around $\phi\sim T_{p}/g_{x}$, enabling the supercooled phase transition. In the vicinity of $\phi\sim T_{p}/g_{x}$, the temperature-dependent contributions can be comparable to the field-dependent terms. We therefore retain all thermal contributions in both the leading and subleading orders of effective potentials. Consequently, the leading and subleading order effective potentials are thus written as
$V^{\rm LT}_{g_{x}^{4}}$ $=-\frac{\mu_{x}^{2}}{2}\phi_{c}^{2}+\frac{\lambda_{x}}{4}\phi_{c}^{4}+3\frac{m_{A^{\prime}}^{4}}{64\pi^{2}}\Bigl(\log\frac{m_{A^{\prime}}^{2}}{\Lambda^{2}}-\frac{5}{6}\Bigr)$
$\quad+V_{{\rm daisy}}^{A^{\prime}}(T^{3})+V_{{\rm daisy}}^{h_{x}}(T^{3})+\frac{T^{4}}{2\pi^{2}}\left[J_{B}\Bigl(\frac{m_{h_{x}}^{2}}{T^{2}}\Bigl)+3J_{B}\Bigl(\frac{m_{A^{\prime}}^{2}}{T^{2}}\Bigl)\right]\,$, (3.93)
$V^{\rm LT}_{g_{x}^{6}}$ $=\frac{m_{G}^{4}}{64\pi^{2}}\Bigl(\log\frac{m_{G}^{2}}{\Lambda^{2}}-\frac{3}{2}\Bigr)-\frac{m_{c}^{4}}{64\pi^{2}}\Bigl(\log\frac{m_{c}^{2}}{\Lambda^{2}}-\frac{3}{2}\Bigr)$
$\quad+V_{{\rm daisy}}^{G}(T^{3})+\frac{T^{4}}{2\pi^{2}}\left[J_{B}\Bigl(\frac{m_{G}^{2}}{T^{2}}\Bigr)-J_{B}\Bigl(\frac{m_{c}^{2}}{T^{2}}\Bigr)\right]\,$, (3.94)
$Z\left(\phi_{c}\right)$ $=1+Z_{g_{x}^{2}}\left(\phi_{c},T=0\right)+\delta Z_{g_{x}^{2}}\left(\phi_{c},T\right)+\mathcal{O}(g_{x}^{4})\,$, (3.95)
where we incorporate the gauge-independent thermal contributions into the leading order effective potential, while the $\xi$-dependent thermal terms are assigned to the subleading effective potential.
The bounce solution is obtained by solving
$\nabla^{2}\phi_{b}=\frac{\partial V^{\rm LT}_{g_{x}^{4}}}{\partial\phi_{c}}\biggr|_{\phi_{c}=\phi_{b}}\,,\qquad\phi_{b}(\infty)=0\,,\qquad\phi_{b}^{\prime}(0)=0\,$, (3.96)
and the nucleation rate is
$\Gamma$ $=\mathcal{A}\,{\rm e}^{-(S^{\rm LT}_{0}+S^{\rm LT}_{1})/T}\,$, (3.97)
$S^{\rm LT}_{0}$ $=\int{\rm d}^{3}x\,\Bigl[V^{\rm LT}_{g_{x}^{4}}(\phi_{b})+\frac{1}{2}(\partial_{i}\phi_{b})^{2}\Bigr]\,$, (3.98)
$S^{\rm LT}_{1}$ $=\int{\rm d}^{3}x\,\Bigl[V^{\rm LT}_{g_{x}^{6}}(\phi_{b})+\frac{1}{2}Z_{g_{x}^{2}}(\partial_{i}\phi_{b})^{2}\Bigr]\,$, (3.99)
where all terms in $S_{0}$ are $\xi$-independent. The Nielsen identities are analogous to the zero-temperature case as
$\xi\frac{\partial V^{\rm LT}_{g_{x}^{4}}}{\partial\xi}$ $=0\,$, (3.100)
$\xi\frac{\partial V^{\rm LT}_{g_{x}^{6}}}{\partial\xi}$ $=-C_{g_{x}^{2}}\frac{\partial V^{\rm LT}_{g_{x}^{4}}}{\partial\phi_{c}}\,$, (3.101)
$\xi\frac{\partial Z_{g_{x}^{2}}}{\partial\xi}$ $=-2\frac{\partial C_{g_{x}^{2}}}{\partial\phi_{c}}\,.$ (3.102)
where $Z_{g_{x}^{2}}$ and $C_{g_{x}^{2}}$ are given in Eqs. (3.42)–(3.43), as in the zero-temperature analysis. Using the identities above, we obtain
$\xi\frac{\partial S^{\rm LT}_{1}}{\partial\xi}$ $=\int{\rm d}^{3}x\,\Bigl[\xi\frac{\partial V^{\rm LT}_{g_{x}^{6}}}{\partial\xi}+\frac{1}{2}\xi\frac{\partial Z_{g_{x}^{2}}}{\partial\xi}\,(\partial_{i}\phi_{b})^{2}\Bigr]$
$=-\int{\rm d}^{3}x\,\Bigl[C_{g_{x}^{2}}\frac{\partial V^{\rm LT}_{g_{x}^{4}}}{\partial\phi_{c}}+\frac{\partial C_{g_{x}^{2}}}{\partial\phi_{c}}\,(\partial_{i}\phi_{b})^{2}\Bigr]\biggr|_{\phi_{c}=\phi_{b}}$
$=-\int{\rm d}^{3}x\,\Bigl[C_{g_{x}^{2}}\frac{\partial V^{\rm LT}_{g_{x}^{4}}}{\partial\phi_{c}}+(\partial_{i}C_{g_{x}^{2}})\,(\partial_{i}\phi_{b})\Bigr]\biggr|_{\phi_{c}=\phi_{b}}$
$=-\int{\rm d}^{3}x\,C_{g_{x}^{2}}\Bigl[\frac{\partial V^{\rm LT}_{g_{x}^{4}}}{\partial\phi_{c}}-\nabla^{2}\phi_{b}\Bigr]\biggr|_{\phi_{c}=\phi_{b}}$
$=0\,$, (3.103)
demonstrating that the subleading effective action $S_{1}$ is gauge invariant.
In practice, once the thermal contributions are included, Eq. (3.101) is still satisfied to a good approximation. The subleading effective action $S^{\rm LT}_{1}$ is shown numerically to be approximately gauge-independent.
3.2.5 Comments and discussions
Regarding gauge-independent treatments of the finite-temperature effective potential and their use in gravitational wave calculations, a few practical points are worth recording:
  • The power counting rule: For the high-temperature phase transition analysis in [39, 43, 44], the Higgs quartic coupling is taken to scale as $\lambda_{x}\sim g_{x}^{3}$. More precisely, it is assumed to lie in the region $g_{x}^{4}\ll\lambda_{x}\ll g_{x}^{3}$ [3]. [^6] For high-temperature phase transitions, one expects $\frac{g_{x}\phi_{c}}{T_{n}}\sim\frac{g_{x}\phi_{c}}{T_{c}}\ll 1\,.$ (3.104) Using the thermal effective potential in Eq. (3.57), one finds at $T_{c}$, $\frac{\phi_{c}}{T_{c}}\sim\frac{2E}{\lambda_{x}}\sim\frac{g_{x}^{3}}{\lambda_{x}}\,.$ (3.105) Substituting this into the previous condition gives $g_{x}\,\frac{\phi_{c}}{T_{c}}\sim\frac{g_{x}^{4}}{\lambda_{x}}\ll 1\,.$ (3.106) Hence, a consistent high-temperature phase transition typically requires $\lambda_{x}\gg g_{x}^{4}$, implying that, for fixed $\lambda_{x}$, moving deeper into the high-temperature regime favors smaller values of $g_{x}$. In our parameter scan, we find that all viable high-temperature phase transitions satisfy $\lambda_{x}\lesssim g_{x}^{3}$, while in many cases $\lambda_{x}\ll g_{x}^{4}$. This typically occurs for benchmark points in which $T_{n}$ lies somewhat below $T_{c}$, while the high-temperature approximation $m_{A^{\prime}}/T_{n}\ll 1$ remains reasonably valid.
    Nevertheless, adopting the scaling $\lambda_{x}\sim g_{x}^{3}$ to organize the leading and subleading terms in the potential does yield a gauge-independent effective action, as can be verified analytically using the Nielsen identity. However, alternative scalings, e.g., $\lambda_{x}\sim g_{x}^{4}$, do not yield a consistent result, as the Nielsen identity is not satisfied at the corresponding order, and thus the resulting subleading effective action is not gauge-independent.
    What, then, is the role of the power counting? In our view, it provides a systematic organization of the calculation: it singles out the leading contributions that determine the bounce solution, treats subleading terms as controlled corrections, and consistently discards higher-order terms in the gauge coupling $g_{x}$.
  • Gauge-parameter dependence: In the gauge-independent treatment, one must ensure that the terms neglected from the full effective potential are indeed parametrically small. Since some of the omitted contributions scale as $\sim\xi^{n}g^{\text{higher powers}}$, the gauge parameter $\xi$ cannot be taken arbitrarily large. Otherwise the power counting scheme ceases to be reliable and the expansion breaks down, even if the Nielsen identity is satisfied up to subleading order in the effective potential. In practice, we find that the gauge-independent treatment remains valid for moderately small values of $\xi$, e.g., $\xi\lesssim\mathcal{O}(1-10)$.
  • Solving for the bounce: A central step in our gravitational wave analysis is to determine the bounce solution associated with the finite-temperature effective potential. In the usual (gauge-dependent) approach, one typically works in the Landau gauge and solves for the bounce using the full effective potential evaluated at $\xi=0$. By contrast, in the gauge-independent treatment one computes the bounce using the leading-order potential $V_{0}$, which is explicitly $\xi$-independent. For this procedure to be self-consistent, the theory must remain in the weak-coupling regime, $\alpha_{x}\ll 1$, such that the subleading contribution $V_{1}$ does not significantly distort the bounce solution determined from $V_{0}$. In practice, we enforce this by restricting the Monte Carlo scan to $g_{x}<0.7$. A more accurate treatment would therefore require solving for the bounce using the full thermal effective potential, which we leave for future work.
  • Comments on the low-temperature analysis: Our scan indicates that a supercooled phase transition typically favors a moderately large gauge coupling, [^7] A conformal $U(1)$ Higgs sector typically contains no tree-level $\phi^{2}$ term. The local minimum at the origin is instead stabilized by the finite-temperature thermal mass. In this setup, a potential barrier can arise even without relying on the thermal cubic contribution $\sim T\,\phi^{3}$, and may be generated predominantly by the Coleman-Weinberg potential. As a result, the allowed range of the gauge coupling $g_{x}$ can be more flexible than in the conventional Higgsed $U(1)$ case with a tachyonic mass term. Moreover, conformal $U(1)$ models naturally admit shallow potentials, and it is therefore common to find $T_{n,p}\ll v$ and $\alpha\gtrsim 1$, corresponding to strong supercooled phase transitions that yield enhanced gravitational wave signals with peak frequencies shifted to lower bands [10, 52, 54, 80, 21, 31, 63]. Hence conformal hidden sectors are often discussed as natural pathways to nanohertz band signals. e.g., $g_{x}\gtrsim 0.5$, in sharp contrast to the high-temperature case, where smaller $g_{x}$ is generally preferred, c.f, Eq. (3.106). This trend can be understood from two main considerations.
    First, as discussed in the previous section, the low-temperature barrier is generated by the thermal cubic term $\sim T\phi^{3}$ near $\phi=0$, whose coefficient scales as $E\propto g_{x}^{3}$. A larger $g_{x}$ therefore enhances the cubic term and makes it easier to develop a sufficiently strong barrier for the supercooled phase transition. [^8] The effective thermal potential in Eq. (3.57) is derived in the high-temperature limit, where the $J_{B}$ functions admit a high-temperature expansion. In the low-temperature regime, Eq. (3.57) remains applicable in the field range $\phi\lesssim T_{p}/g_{x}$. In particular, as $\phi\to 0$ the field value is again much smaller than the temperature, and thus the high-temperature expansion remains valid in this small-field region, which is the most relevant for generating a thermal barrier. Nevertheless, Eq. (3.57) should not be used indiscriminately across the full temperature range in analyses of supercooled phase transitions.
    Second, the tree-level tachyonic mass term must be lifted to an effective positive value by thermal mass corrections proportional to $g_{x}^{2}$ in order to create and stabilize a local minimum at $\phi=0$. A larger $g_{x}$ thus keeps the symmetric point $\phi=0$ metastable down to lower temperatures, delaying the loss of metastability, which is essential for realizing a supercooled transition.
4 Gravitational waves from first-order phase transition
In a $U(1)$ gauge theory, the finite-temperature effective potential can induce a first-order phase transition and consequently generate a stochastic gravitational wave background. In this section, we review the standard computation of the gravitational wave signal, starting from the nucleation rate inferred from the bounce solution for a given set of model parameters.
4.1 The tunneling action and transition rate
As the temperature falls below the critical temperature $T_{c}$, the effective potential may develop two local minima separated by a barrier, rendering the minimum in the symmetric phase metastable. The transition from the false vacuum to the true vacuum then proceeds via bubble nucleation. The resulting bubbles expand and collide, sourcing a stochastic gravitational wave background. At finite temperature, the bubble nucleation rate, also referred to as the transition rate or the false vacuum decay rate, is given by [78, 79]
$\Gamma(T)=\mathcal{A}(T)\,{\rm e}^{-S_{3}(T)/T}\,$, (4.1)
where $S_{3}(T)\equiv S_{3}(\phi_{b},T)$ is the three-dimensional Euclidean action evaluated on the $O(3)$ symmetric bounce configuration. The action is defined as
$S_{3}(\phi,T)=\int_{0}^{\infty}{\rm d}^{3}x\,\bigl[\tfrac{1}{2}(\nabla\phi)^{2}+\widetilde{V}_{{\rm eff}}(\phi,T)\bigr]\,$, (4.2)
with $\widetilde{V}_{{\rm eff}}(\phi,T)\equiv V_{{\rm eff}}(\phi,T)-V_{{\rm eff}}(\phi_{{\rm FV}},T)$ the free energy density shifted relative to the false vacuum. The bounce $\phi_{b}$ satisfies the equation of motion
$\frac{{\rm d}^{2}\phi}{{\rm d}r^{2}}+\frac{2}{r}\frac{{\rm d}\phi}{{\rm d}r}=\frac{\partial}{\partial\phi}\widetilde{V}_{{\rm eff}}(\phi,T)\,$, (4.3)
and subjects to the boundary conditions $\phi(r\to\infty)=0$ and ${\rm d}\phi/{\rm d}r\bigl|_{r=0}=0$, with $r^{2}=\bm{x}^{2}$. In this work the solution is obtained using CosmoTransitions [81].
The prefactor $\mathcal{A}(T)$ is given by [82]
$\mathcal{A}(T)=T\biggl[\frac{S_{3}(\phi_{b},T)}{2\pi T}\biggr]^{3/2}\Biggl\{\frac{{\rm det}^{\prime}\bigl[-\nabla^{2}+V^{\prime\prime}(\phi_{b},T)\bigr]}{{\rm det}\bigl[-\nabla^{2}+V^{\prime\prime}(\phi_{{\rm FV}},T)\bigr]}\Biggr\}^{-1/2}\,$, (4.4)
where $\rm det^{\prime}$ denotes the functional determinant with zero eigenvalues omitted. For dimensional consistency, $\mathcal{A}(T)$ is commonly approximated as [82]
$\mathcal{A}(T)\simeq T^{4}\biggl[\frac{S_{3}(\phi_{b},T)}{2\pi T}\biggr]^{3/2}\,.$ (4.5)
The false vacuum fraction $P_{f}$ is introduced to track the progress of the phase transition, written as [83, 84]
$P_{f}(t)=\exp\bigl[-\mathcal{V}_{t}^{{\rm ext}}(t)\bigr]\,$, (4.6)
where $\mathcal{V}_{t}^{{\rm ext}}(T)$ denotes the fractional extended volume occupied by true vacuum bubbles at time $t$, given by
$\mathcal{V}_{t}^{{\rm ext}}(t)=\frac{4\pi}{3}\int_{t_{c}}^{t}{\rm d}t^{\prime}\,\Gamma(t^{\prime})\frac{a^{3}(t^{\prime})}{a^{3}(t)}R^{3}(t^{\prime},t)\,$, (4.7)
with
$R(t^{\prime},t)=\int_{t^{\prime}}^{t}{\rm d}t^{\prime\prime}\,v_{w}\frac{a(t)}{a(t^{\prime\prime})}\,$, (4.8)
in which an initial bubble radius is often neglected and the bubble wall velocity $v_{w}$ is assumed to remain constant throughout the expansion. Thus Eq. (4.7) becomes
$\mathcal{V}_{t}^{{\rm ext}}(t)=\frac{4\pi}{3}v_{w}^{3}\int_{t_{c}}^{t}{\rm d}t^{\prime}\,\Gamma(t^{\prime})\biggl[\frac{a(t^{\prime})}{a(t)}\biggr]^{3}\biggl[\int_{t^{\prime}}^{t}{\rm d}t^{\prime\prime}\frac{a(t)}{a(t^{\prime\prime})}\biggr]^{3}\,.$ (4.9)
4.2 Characteristic temperatures
To analyze the transition dynamics and predict the resulting gravitational waves, we introduce several benchmark temperatures that characterize key stages of the phase transition.
Critical temperature $T_{c}$
The critical temperature $T_{c}$ is defined as the temperature at which the effective potential of the true and false vacua become degenerate, $V_{{\rm eff}}(\phi_{{\rm TV}},T_{c})=V_{{\rm eff}}(\phi_{{\rm FV}},T_{c})$, with a potential barrier separating the two minima.
Nucleation temperature $T_{n}$
The nucleation temperature $T_{n}$ is the temperature at which bubbles of the broken phase begin to form in the false vacuum at an appreciable rate, signaling the onset of efficient nucleation. It is defined when one bubble is nucleated within one Hubble volume per Hubble time
$\int_{t_{c}}^{t_{n}}\frac{\Gamma}{H^{3}}\,{\rm d}t\simeq 1\,$, (4.10)
where $\Gamma$ is the bubble nucleation rate and $H$ is the Hubble parameter including the vacuum energy contribution, given by
$H^{2}=\frac{\rho_{{\rm rad}}+\rho_{{\rm vac}}}{3M_{{\rm Pl}}^{2}}=\frac{1}{3M_{{\rm Pl}}^{2}}\Bigl[\frac{\pi^{2}}{30}g_{\ast}T^{4}+\Delta V_{{\rm eff}}(T)\Bigr]\,$, (4.11)
where $\Delta V_{{\rm eff}}=V_{{\rm eff}}(\phi_{{\rm FV}},T)-V_{{\rm eff}}(\phi_{{\rm TV}},T)$, $g_{\ast}$ is the effective number of relativistic degrees of freedom including both the SM and hidden sectors, and $M_{{\rm Pl}}=2.4\times 10^{18}$ GeV denotes the reduced Planck mass.
Given the temperature dependence of both $\Gamma(T)$ and $H(T)$, it is convenient to express the nucleation condition as a function of temperature [85]. Assuming an adiabatically expanding Universe, the conservation of entropy per comoving volume implies
$\frac{{\rm d}}{{\rm d}t}\bigl[s(t)a^{3}(t)\bigr]=0\;\;\Rightarrow\;\;\frac{{\rm d}s}{{\rm d}t}=-3H(t)s(t)\,$, (4.12)
where $s$ is the entropy density of the plasma and $a$ is the scale factor. The entropy density is written as
$s(T)=\frac{\partial p}{\partial T}=-\frac{\partial V}{\partial T}\,$, (4.13)
which can be combined with Eq. (4.12) and obtain
$\frac{{\rm d}T}{{\rm d}t}\frac{{\rm d}s}{{\rm d}T}=-\frac{{\rm d}T}{{\rm d}t}\frac{\partial^{2}V}{\partial T^{2}}=3H(T)\frac{\partial V}{\partial T}\,$, (4.14)
yielding
$\frac{{\rm d}T}{{\rm d}t}=-3H(T)\frac{\partial_{T}V}{\partial_{TT}V}\,$, (4.15)
where the derivatives $\partial_{T}V=\frac{\partial V}{\partial T}$ and $\partial_{TT}V=\frac{\partial^{2}V}{\partial T^{2}}$ are evaluated at the false vacuum. In the literature, the MIT bag equation of state [86] is often employed, in which the false vacuum potential is parameterized as
$V(\phi_{{\rm FV}},T)=aT^{4}+b\,$, (4.16)
with $a,b$ being temperature-independent coefficients. This bag model simplifies Eq. (4.15) to the familiar form
$\frac{{\rm d}T}{{\rm d}t}\stackrel{{\scriptstyle{\rm bag}}}{{=}}-H(T)T\,$, (4.17)
and the nucleation condition defined by Eq. (4.10) can then be rewritten as
$\int_{T_{n}}^{T_{c}}\frac{\Gamma(T)}{H(T)^{4}}\frac{1}{T}\,{\rm d}T\simeq 1\,.$ (4.18)
When the phase transition occurs near the electroweak scale, the nucleation condition in Eq. (4.18) can be well approximated by $S_{3}(T_{n})/T_{n}\simeq 140$.
For a high-temperature phase transition, the nucleation temperature $T_{n}$ typically lies only slightly below $T_{c}$ and marks the onset of the transition. This is not the case for a supercooled phase transition, where the onset is instead characterized by the percolation temperature $T_{p}$.
Percolation temperature $T_{p}$
The percolation temperature $T_{p}$ is the temperature at which bubbles have nucleated and grown enough that they percolate through space.
For a fast phase transition, the nucleation temperature $T_{n}$ is sufficient to characterize both the onset and completion of the transition. However, a slow or supercooled transition is long lasting, during which the Universe can remain trapped in the false vacuum for an extended period after it becomes metastable, delaying bubble collision to the temperature much lower than $T_{n}$. In such case, $T_{p}$ provides a more accurate description of the transition dynamics and is supposed as the reference temperature for gravitational wave calculations, particularly for the strong supercooling in which $T_{p}\ll T_{n}$ [50].
The percolation temperature $T_{p}$ is defined by roughly $70\%$ of the Universe remains in the false vacuum, $P_{f}(T_{p})\approx 0.7$ [87, 48]. In Eq. (4.9), the ratio of scale factors at different times is given by
$\frac{a(t_{1})}{a(t_{2})}=\exp\Bigl[\int_{t_{2}}^{t_{1}}{\rm d}t^{\prime}H(t^{\prime})\Bigr]\,.$ (4.19)
Assuming an adiabatic expansion and adopting the MIT bag equation of state, this ratio can be recast as a function of temperature
$\frac{a(T_{1})}{a(T_{2})}=\exp\Bigl(\int_{T_{1}}^{T_{2}}{\rm d}T^{\prime}\frac{1}{T^{\prime}}\Bigr)=\frac{T_{2}}{T_{1}}\,$, (4.20)
leading to
$P_{f}(T)=\exp\biggl\{-\frac{4\pi}{3}v_{w}^{3}\int_{T}^{T_{c}}{\rm d}T^{\prime}\frac{\Gamma(T^{\prime})}{T^{\prime 4}H(T^{\prime})}\Bigl[\int_{T}^{T^{\prime}}\frac{{\rm d}T^{\prime\prime}}{H(T^{\prime\prime})}\Bigr]^{3}\biggr\}\,.$ (4.21)
To ensure the phase transition completes successfully, two conditions must be checked [85, 60]: (1) the transition has an end temperature $T_{\rm e}$, defined by $P_{f}(T_{\rm e})=\varepsilon\lesssim 0.01$; (2) the physical volume of the false vacuum $\mathcal{V}_{{\rm phys}}(t)=a^{3}(t)P_{f}(t)$ decreases with time,
$\frac{{\rm d}\mathcal{V}_{{\rm phys}}}{{\rm d}t}=\mathcal{V}_{{\rm phys}}\Bigl[\frac{{\rm d}}{{\rm d}t}\ln P_{f}(t)+3H(t)\Bigr]\leq 0\,$, (4.22)
giving rise to
$T\frac{{\rm d}\mathcal{V}_{t}^{{\rm ext}}}{{\rm d}T}+3\leq 0\,$, (4.23)
and this condition should be verified at both $T_{p}$ and $T_{\rm e}$.
Reheating temperature $T_{\rm reh}$
A first-order phase transition releases latent heat and can reheat the plasma from the percolation temperature $T_{p}$ to a reheating temperature $T_{\rm reh}$. Assuming instantaneous reheating, $T_{{\rm reh}}$ is estimated by
$T_{{\rm reh}}\simeq\bigl[1+\alpha(T_{p})\bigr]^{1/4}T_{p}\,$, (4.24)
where $\alpha(T)$ is the transition strength at temperature $T$. For moderate supercooling ($\alpha\lesssim 1$), one has $T_{\rm reh}\approx T_{p}$; whereas for strong supercooling ($\alpha\gg 1$), the reheating temperature can be significantly higher than $T_{p}$, or even higher than $T_{c}$ [60]. The present gravitational wave amplitude and frequency are obtained by redshifting their values from the reheating temperature $T_{\rm reh}$ rather than from $T_{p}$ down to today.
To summarize this subsection, as the Universe cools, the dark $U(1)_{x}$ sector undergoing a supercooled phase transition may pass through the following characteristic temperatures,
$T_{c}>T_{n}\gg T_{p}\gtrsim T_{\rm e}\gtrsim T_{0}\gtrsim 0\,.$ (4.25)
Recall that $T_{0}$ is defined as the temperature at which $\phi=0$ ceases to be a local minimum. In a typical supercooled phase transition, $T_{p}$, $T_{\rm e}$, and $T_{0}$ are close to one another.
4.3 Hydrodynamic parameters
The gravitational wave signal originating from a first-order phase transition depends sensitively on the characteristic length scale $L_{\ast}$, which represents the typical scale carrying most of the released energy. A common estimate identifies $L_{\ast}$ with the mean bubble separation $R_{\ast}$, which is
$R_{\ast}=\frac{(8\pi)^{1/3}}{\beta}v_{w}\,$, (4.26)
where $v_{w}$ is the bubble wall velocity and $\beta$ is the inverse time scale, defined as
$\beta=-\frac{{\rm d}}{{\rm d}t}\Bigl(\frac{S_{3}}{T}\Bigr)\biggr|_{t=t_{p}}=H(T)T\frac{{\rm d}}{{\rm d}T}\Bigl(\frac{S_{3}}{T}\Bigr)\biggr|_{T=T_{p}}\,$, (4.27)
which corresponds to an exponential nucleation rate and is usually expressed in the following form
$\frac{\beta}{H_{p}}=T\frac{{\rm d}}{{\rm d}T}\Bigl(\frac{S_{3}}{T}\Bigr)\biggr|_{T=T_{p}}\,.$ (4.28)
For strong supercooled phase transitions, $\beta$ may become tiny or even negative, indicating the commonly exponential approximation breaks down. In this case, the mean bubble separation $R_{\ast}$ should instead be directly determined from the bubble number density
$R_{\ast}(T)\equiv\bigl[n(T)\bigr]^{-1/3}=\biggl[T\int_{T}^{T_{c}}{\rm d}T^{\prime}\frac{\Gamma(T^{\prime})P_{f}(T^{\prime})}{T^{\prime 4}H(T^{\prime})}\biggr]^{-1/3}\,.$ (4.29)
The strength of the phase transition is parameterized by $\alpha$, defined as [88, 89]
$\alpha\equiv\frac{\bar{\theta}_{f}(T_{p})-\bar{\theta}_{t}(T_{p})}{3w_{f}(T_{p})}\,,\qquad\bar{\theta}\equiv e-\frac{p}{c_{s,t}^{2}}\,$, (4.30)
where $w_{f}$ denotes the enthalpy density in the false vacuum, $\bar{\theta}$ represents the pseudotrace that is a generalization of the vacuum energy, and $c_{s}^{2}$ is the speed of sound in the plasma. Subscripts $f$ and $t$ refer to quantities in the false and true vacua respectively. The hydrodynamics of the plasma is described by the energy density $e$, pressure $p$, and enthalpy density $w$. The plasma is usually modeled as a perfect relativistic fluid, for which the pressure is related to the free energy density $\mathcal{F}$ by $p=$-$\mathcal{F}$. Since $\mathcal{F}$ equals to the effective potential, one has $p=-V_{{\rm eff}}$. The relevant thermodynamic quantities are given by
$p$ $\equiv-V_{{\rm eff}}(\phi_{c},T)+\frac{\pi^{2}}{90}g_{{\rm eff}}(T)T^{4}\,$, (4.31)
$e$ $\equiv T\frac{\partial p}{\partial T}-p\,$, (4.32)
$w$ $\equiv T\frac{\partial p}{\partial T}=p+e\,$, (4.33)
$c_{s}^{2}$ $\equiv\frac{\delta p}{\delta e}\simeq\frac{\partial p/\partial T}{\partial e/\partial T}\,$, (4.34)
where the second term in $p$ accounts for the field-independent contribution from all relativistic particle species in thermal equilibrium with the dark Higgs $h_{x}$ at temperature $T$ [89].
During the phase transition, the released energy accelerates bubbles, while the interaction between bubble walls and the surrounding plasma impedes the acceleration. When the released energy is sufficiently large to overcome this friction, the wall continues to accelerate without reaching a terminal velocity, entering a runaway regime. The corresponding criterion parameter is defined as [90]
$\alpha_{\infty}\equiv\frac{1}{18}\frac{\sum_{i}g_{i}c_{i}\Delta m_{i}^{2}T_{p}^{2}}{w_{f}(T_{p})}\,$, (4.35)
where $c_{i}=1\,(1/2)$ for bosons (fermions), $g_{i}$ denotes the internal degrees of freedom of particle $i$ and $\Delta m_{i}^{2}=m_{i,t}^{2}-m_{i,f}^{2}$ with the sum running over all species that gain a mass during the transition. If $\alpha>\alpha_{\infty}$, the bubble wall runs away, yielding $v_{w}\simeq 1$. In the non-runaway regime ($\alpha<\alpha_{\infty}$), the wall eventually reaches a model-dependent terminal velocity. We compute the wall velocity $v_{w}(\alpha,c_{s,f}^{2},c_{s,t}^{2},\Psi)$, where $\Psi\equiv w_{t}/w_{f}$ denotes the ratio of the enthalpy densities in the true and false phases, following [91, 92].
Kinetic energy fraction is defined as $K\equiv\rho_{{\rm fl}}/e_{p}$ which quantifies the fraction of available vacuum energy converted into bulk motion of the plasma, where $e_{p}$ is the total energy density at the percolation temperature $T_{p}$, and $\rho_{{\rm fl}}$ is the plasma kinetic energy density, obtained by integrating over the fluid profile. The kinetic energy fraction can be further parameterized as
$K\simeq\frac{\alpha}{1+\alpha}\kappa(\alpha,c_{s,f}^{2},c_{s,t}^{2},v_{w})\,$, (4.36)
where $\kappa$ represents the efficiency factor, determined by the transition strength, sound speeds in two phases and wall velocity, with detailed calculations provided in [88, 89].
4.4 Gravitational wave power spectrum
The gravitational wave power spectrum is defined as
$\Omega_{{\rm GW}}(f)\equiv\frac{1}{\rho_{c}}\frac{{\rm d}\rho_{{\rm GW}}}{{\rm d}\log f}\,$, (4.37)
where $f$ denotes the frequency and $\rho_{c}$ is the critical density. The total gravitational wave signal typically receives three primary contributions: (1) bubble collisions of the scalar field; (2) sound waves in the bulk plasma; (3) magneto-hydrodynamic turbulence in the plasma [93, 94, 95]. Accordingly,
$h^{2}\Omega_{{\rm GW}}(f)=h^{2}\Omega_{\phi}(f)+h^{2}\Omega_{{\rm sw}}(f)+h^{2}\Omega_{{\rm tb}}(f)\,$, (4.38)
with each component parametrized as
$h^{2}\Omega(f)=h^{2}\Omega^{{\rm peak}}\mathcal{S}(f)\,$, (4.39)
where $\Omega^{{\rm peak}}$ is the peak amplitude and $\mathcal{S}$ the spectral-shape function.
Phase transitions in a dark sector can be characterized by two strength parameters, $\alpha_{{\rm tot}}$ and $\alpha_{h}$, defined as [17, 22, 25, 33],
$\alpha_{{\rm tot}}=\frac{\Delta\bar{\theta}(T_{h,p})}{3\bigl[w_{f}^{v}(T_{v,p})+w_{f}^{h}(T_{h,p})\bigr]}\,,\qquad\alpha_{h}=\frac{\Delta\bar{\theta}(T_{h,p})}{3w_{f}^{h}(T_{h,p})}\,$, (4.40)
where $\Delta\bar{\theta}=\bar{\theta}_{f}-\bar{\theta}_{t}$, c.f., Eq. (4.30). Here, $w_{f}^{v}$ and $w_{f}^{h}$ denote the false vacuum enthalpy densities of the visible and hidden sector, respectively, and $T_{v,p}$ and $T_{h,p}$ are their corresponding percolation temperatures. In this work, we assume that the dark sector shares the same temperature as the visible sector, $T_{h}=T_{v}$, due to efficient interactions between the two sectors. The total parameter $\alpha_{{\rm tot}}$ determines the amplitude of gravitational waves, and $\alpha_{h}$ controls the efficiency factor $\kappa$ which describes the fraction of vacuum energy transformed into the bulk kinetic energy of the plasma.
Under this framework, the runaway condition introduced in Eq. (4.35) generalizes to the hidden sector as
$\alpha_{h,\infty}\equiv\frac{1}{18}\frac{\sum_{i}g_{i}c_{i}\Delta m_{i}^{2}T_{h,p}^{2}}{w_{f}^{h}(T_{h,p})}\,$, (4.41)
where $c_{i}=1\,(1/2)$ for bosons (fermions), $g_{i}$ is the internal degrees of freedom of particle $i$, and $\Delta m_{i}^{2}=m_{i,t}^{2}-m_{i,f}^{2}$. In the runaway regime ($\alpha_{h,\infty}<\alpha_{h}$), efficiency factors are given by
$\kappa_{{\rm col}}=1-\frac{\alpha_{h,\infty}}{\alpha_{h}}\,,\quad\kappa_{{\rm sw}}=\frac{\alpha_{h,\infty}}{\alpha_{h}}\,\kappa(\alpha_{h,\infty},c_{s,f}^{2},c_{s,t}^{2},v_{w})\,,\quad\kappa_{{\rm tb}}=\epsilon\kappa_{{\rm sw}}\,$, (4.42)
where $\epsilon$ denotes the fraction of bulk motion that is turbulent. In the non-runaway regime ($\alpha_{h,\infty}>\alpha_{h}$), the factors are
$\kappa_{{\rm col}}=0\,,\quad\kappa_{{\rm sw}}=\kappa(\alpha_{h},c_{s,f}^{2},c_{s,t}^{2},v_{w})\,,\quad\kappa_{{\rm tb}}=\epsilon\kappa_{{\rm sw}}\,$, (4.43)
In this work we adopt $\epsilon=0.1$ according to numerical simulations [96].
The present gravitational wave amplitude and frequency are obtained by redshifting their values at production evaluated at the temperature $T_{\ast}$, given by
$h^{2}\Omega_{0}(f_{0})=\Bigl(\frac{a_{\ast}}{a_{0}}\Bigr)^{4}\Bigl(\frac{H_{\ast}}{H_{0}}\Bigr)^{2}\Omega_{1}(f_{\ast})=1.67\times 10^{-5}\biggl(\frac{100}{g_{{\rm eff}}(T_{\ast})}\biggr)^{1/3}\Omega_{\ast}(f_{\ast})\,$, (4.44)
and
$f_{0}=\frac{a_{\ast}}{a_{0}}f_{\ast}=1.65\times 10^{-5}{\rm Hz}\,\biggl(\frac{g_{{\rm eff}}(T_{\ast})}{100}\biggr)^{1/6}\biggl(\frac{T_{\ast}}{100\,{\rm GeV}}\biggr)\,$, (4.45)
where the subscript 0 denotes present quantities and $\ast$ denotes those evaluated at the time of the gravitational wave production.
As noted in the previous section, a supercooled phase transition releases latent heat and can reheat the plasma from the percolation temperature $T_{p}$ to a reheating temperature $T_{\rm reh}$. The present gravitational wave amplitude and frequency are therefore obtained by evaluating the redshift factor at $T_{\ast}=T_{\rm reh}$ rather than at $T_{p}$.
The gravitational wave power spectra observed today for the three contributions are thus given by
  1. 1.
    bubble collision
    $h^{2}\Omega_{\phi}(f)$ $=1.67\times 10^{-5}\frac{0.11v_{w}^{3}}{0.42+v_{w}^{2}}\Bigl(\frac{100}{g_{{\rm eff}}(T_{{\rm reh}})}\Bigr)^{1/3}\Bigl(\frac{\kappa_{{\rm col}}\alpha_{{\rm tot}}}{1+\alpha_{{\rm tot}}}\Bigr)^{2}\Bigl(\frac{\beta}{H_{\ast}}\Bigr)^{-2}\mathcal{S}_{\phi}(f)\,$, (4.46)
    $\mathcal{S}_{\phi}(f)$ $=\frac{3.8(f/f_{\phi})^{2.8}}{1+2.8(f/f_{\phi})^{3.8}}\,$, (4.47)
    $f_{\phi}$ $=1.6\times 10^{-7}\,\Bigl(\frac{g_{{\rm eff}}(T_{{\rm reh}})}{100}\Bigr)^{1/6}\Bigl(\frac{T_{{\rm reh}}}{1\,{\rm GeV}}\Bigr)\Bigl(\frac{\beta}{H_{\ast}}\Bigr)\Bigl(\frac{0.62}{1.8-0.1v_{w}+v_{w}^{2}}\Bigr)\,{\rm Hz}\,.$ (4.48)
  2. 2.
    sound wave
    $h^{2}\Omega_{{\rm sw}}(f)$ $=2.65\times 10^{-6}v_{w}\Bigl(\frac{100}{g_{{\rm eff}}(T_{{\rm reh}})}\Bigr)^{1/3}\Bigl(\frac{\kappa_{{\rm sw}}\alpha_{{\rm tot}}}{1+\alpha_{{\rm tot}}}\Bigr)^{2}\Bigl(\frac{\beta}{H_{\ast}}\Bigr)^{-1}\mathcal{S}_{{\rm sw}}(f)\,$, (4.49)
    $\mathcal{S}_{{\rm sw}}(f)$ $=\Bigl(\frac{f}{f_{{\rm sw}}}\Bigr)^{3}\biggl[\frac{7}{4+3\bigl(f/f_{{\rm sw}}\bigr)^{2}}\biggr]^{7/2}\,$, (4.50)
    $f_{{\rm sw}}$ $=1.9\times 10^{-7}\,\frac{1}{v_{w}}\Bigl(\frac{g_{{\rm eff}}(T_{{\rm reh}})}{100}\Bigr)^{1/6}\Bigl(\frac{T_{{\rm reh}}}{1\,{\rm GeV}}\Bigr)\Bigl(\frac{\beta}{H_{\ast}}\Bigr)\,{\rm Hz}\,.$ (4.51)
  3. 3.
    turbulence
    $h^{2}\Omega_{{\rm tb}}(f)$ $=3.35\times 10^{-4}v_{w}\Bigl(\frac{100}{g_{{\rm eff}}(T_{{\rm reh}})}\Bigr)^{1/3}\Bigl(\frac{\kappa_{{\rm tb}}\alpha_{{\rm tot}}}{1+\alpha_{{\rm tot}}}\Bigr)^{3/2}\mathcal{S}_{{\rm tb}}(f)\,$, (4.52)
    $\mathcal{S}_{{\rm tb}}(f)$ $=\biggl(\frac{f}{f_{{\rm tb}}}\biggr)^{3}\biggl[\frac{1}{1+\bigl(f/f_{{\rm tb}}\bigr)}\biggr]^{11/3}\Bigl(1+8\pi\frac{f}{H_{\ast}^{\prime}}\Bigr)^{-1}\,$, (4.53)
    $f_{{\rm tb}}$ $=\frac{2.7}{1.9}f_{{\rm sw}}\,$, (4.54)
where $f_{\phi},f_{{\rm sw}},f_{{\rm tb}}$ are peak frequencies, $H_{\ast}$ denotes the Hubble rate at gravitational wave production, where we take $T_{\ast}=T_{{\rm reh}}$, and $H_{\ast}^{\prime}$ is the redshifted Hubble rate, given by
$H_{\ast}^{\prime}=\Bigl(\frac{a}{a_{0}}\Bigr)H_{\ast}\simeq 1.6\times 10^{-5}\,{\rm Hz}\,\Bigl(\frac{g_{{\rm eff}}(T_{{\rm reh}})}{100}\Bigr)^{1/6}\Bigl(\frac{T_{{\rm reh}}}{100\,{\rm GeV}}\Bigr)\,.$ (4.55)
Strong signals from supercooling
Using the hydrodynamic parameters defined in Section 4.3 together with the gravitational wave power spectrum formulas, we summarize the main factors by which supercooled phase transitions can generically produce stronger signals:
  • From Eq. (4.30), the transition strength parameter $\alpha$ scales inversely with the radiation energy density, $\rho_{\rm rad}(T_{p})\propto T_{p}^{4}$. Supercooled transitions typically have a much lower percolation temperature $T_{p}$, and thus lead to a large $\alpha$. In this regime the fraction of energy available to source gravitational waves, $\kappa\alpha/(1+\alpha)$, is enhanced and approaches $\kappa$. As a result, the released vacuum energy can constitute a sizable fraction of the total energy density, leading to substantially stronger gravitational wave signals.
  • The parameter $\beta$ characterizes the inverse time scale over which the nucleation rate grows, and hence $\beta/H$ controls the typical duration of the transition. In supercooled scenarios, nucleation is delayed and then turns on rapidly, thus the transition proceeds with larger characteristic bubble sizes; c.f., Eq. (4.26). Since gravitational waves are sourced by anisotropic stresses on length scales of order $R_{\ast}\sim v_{w}/\beta$, a smaller $\beta/H$ typically enhances the signal amplitude.
  • In the strong supercooling limit, the driving pressure can be large while plasma friction becomes relatively less important. Bubble walls may therefore become highly relativistic and can more readily approach runaway behavior. This increases the fraction of the released energy stored in the bubble walls or in the scalar field kinetic energy, thereby enhancing the collision- and field-sourced contributions to the gravitational wave spectrum.
Finally, noted from gravitational wave power spectrum formulas, the peak frequency parametrically scales as
$f_{\rm peak}\,\propto\,T_{\ast}\times\frac{\beta}{H_{\ast}}\times\big(\text{redshift factors}\big)\,.$ (4.56)
Supercooling typically implies a smaller characteristic temperature $T_{\ast}$ and often a smaller $\beta/H_{\ast}$, both of which push $f_{\rm peak}$ to lower values. Consequently, supercooled transitions are a well-motivated and frequently studied mechanism for generating signals in the nanohertz band.
4.5 Gauge-dependent versus gauge-independent results
In this subsection, we perform a Monte Carlo scan of the parameter space for gravitational waves sourced by the first-order phase transition, using both a gauge-dependent treatment in Landau gauge ($\xi=0$) and the gauge-independent framework discussed in Section 3.
4.5.1 Gauge-dependent gravitational wave results
To begin, we perform a scan using the gauge-dependent finite-temperature effective potential in Landau gauge ($\xi=0$), as presented in Section 3.1. We sample $3\times 10^{4}$ parameter points in the ranges
$g_{x}\in[10^{-2},1],\;\lambda_{x}\in[10^{-6},10^{-1}],\;v_{x}\in[10^{-3},10^{3}]\,{\rm GeV}\,.$ (4.57)
and find $6962$ points for which the phase transition completes successfully. The corresponding peak signals are shown in Fig. 1. The scatter points are color-coded by $T_{p}/m_{A^{\prime}}$ and by $v_{x}$.
[…]
Although results obtained from a gauge-dependent potential do not constitute definitive gravitational wave predictions for the $U(1)$ dark sector, the two panels nevertheless reveal clear trends that are instructive for understanding the underlying physics: larger symmetry breaking scales $v_{x}$ shift the gravitational wave signal toward higher frequencies, while smaller $v_{x}$ shift it toward lower frequencies. In our scan, the lowest peak frequencies approach the nanohertz band with $v_{x}\lesssim 100$ MeV.
The temperature at which the transition occurs is also closely correlated with the signal strength: low-temperature transitions typically yield larger amplitudes, whereas transitions occurring at higher temperatures tend to produce much weaker signals.
To quantify the range of thermal regimes encountered in the gauge-dependent scan, we classify the successful transition points into four categories according to the ratio $T_{p}/m_{A^{\prime}}$, with the supercooled and (approximately) high-temperature regimes being the ones that can be treated within the gauge-independent framework:
  1. 1.
    Supercooled phase transition: $T_{p}/m_{A^{\prime}}\lesssim 0.2$;
  2. 2.
    (Moderately) low-temperature phase transition: $0.2<T_{p}/m_{A^{\prime}}<0.5$;
  3. 3.
    Intermediate regime: $0.5<T_{p}/m_{A^{\prime}}<2$;
  4. 4.
    (Approximately) high-temperature phase transition: $T_{p}/m_{A^{\prime}}\gtrsim 2$.
The number of points and the corresponding fractions in each category are summarized in Table 1, and the overall composition is illustrated by the pie chart in Fig. 2.
Case number fraction
Supercooled: $T_{p}/m_{A^{\prime}}\lesssim 0.2$ 2156 30.97 %
(Approx) low-$T$: $0.2<T_{p}/m_{A^{\prime}}<0.5$ 2083 29.92 %
Intermediate: $0.5<T_{p}/m_{A^{\prime}}<2$ 2377 34.14 %
(Approx) high-$T$: $T_{p}/m_{A^{\prime}}\gtrsim 2$ 346 4.97 %
Total 6962 100 %
Table 1: Statistics of successful first-order phase transition points obtained in the gauge-dependent calculation (Landau gauge, $\xi=0$). We sample $3\times 10^{4}$ parameter points in the ranges of Eq. (4.57) and find $6962$ points for which the transition percolates successfully. The successful points are classified by the ratio $T_{p}/m_{A^{\prime}}$. For each category we report the number of points and the corresponding fraction of the total sample.
Figure 2: Pie chart showing the composition of phase transition regimes in the gauge-dependent scan (Landau gauge, $\xi=0$) over the parameter ranges in Eq. (4.57). The classification is based on the ratio $T_{p}/m_{A^{\prime}}$ as defined in Table 1, which also lists the corresponding event counts and fractions.
From the statistics, we find that supercooled transitions account for roughly $30\%$ of all successful phase transition points. They typically yield the strongest gravitational wave signals among the regimes considered, as shown in Fig. 1, and many lie within the projected sensitivity of near-future experiments. Moderately low-temperature transitions with $0.2<T_{p}/m_{A^{\prime}}<0.5$ constitute another $\sim 30\%$ of the successful points. Taken together, the supercooled and moderately low-temperature regimes are naturally and approximately, respectively, suited to the low-temperature gauge-independent framework.
We identify an approximately high-temperature regime using the criterion $T_{p}/m_{A^{\prime}}\gtrsim 2$, which already constitutes a small subset of the successful points (below $5\%$). Truly high-temperature transitions for which the high-temperature expansion is parametrically justified, $T_{p}/m_{A^{\prime}}\gg 10$, are rare in our scan. These points generally produce very weak signals and are therefore difficult to detect, making them phenomenologically less compelling.
The remaining points lie in an intermediate regime, $0.5<T_{p}/m_{A^{\prime}}<2$, and make up about $35\%$ of the successful sample. Their predicted signals are intermediate between the supercooled and high-temperature cases. This regime is not applicable to the gauge-independent treatments: the high-temperature expansion is not valid, while the low-temperature thermal contributions cannot be regarded as small corrections. Moreover, the associated gravitational wave signals are typically below the reach of near-future detectors, rendering this region also less phenomenologically relevant at present.
It should be noted that part of the literature adopts Eq. (3.57) as the starting point for gravitational wave analyses and reports many viable points in the intermediate- and low-temperature regimes. Such treatments are not appropriate. Eq. (3.57) is obtained by employing the high-temperature expansion of the $J_{B}$ functions, which ceases to be valid once the high-temperature condition is no longer satisfied. Consequently, the resulting gravitational wave spectra in regimes where the high-temperature expansion is not valid are unreliable, regardless of gauge-dependent issues.
4.5.2 Gauge-independent gravitational wave predictions
The gauge-independent effective action in the high- and low-temperature limits is discussed in detail in Section 3, and we scan these two regimes separately. The ranges of input parameters are chosen as follows:
$\text{high-}T:\quad$ $g_{x}\in[10^{-3},0.7],\;\lambda_{x}\in[10^{-9},10^{-2}],\;v_{x}\in[10^{-3},10^{2}]\,{\rm GeV}\,$, (4.58)
$\text{low-}T:\quad$ $g_{x}\in[10^{-2},0.7],\;\lambda_{x}\in[10^{-6},10^{-2}],\;v_{x}\in[10^{-3},10^{2}]\,{\rm GeV}\,.$ (4.59)
For each regime, we sample $3\times 10^{4}$ points and compute the resulting stochastic gravitational wave background. The results are displayed in the $\{f,\,\Omega_{\rm GW}h^{2}\}$ plane, where each scatter point corresponds to the peak frequency and peak amplitude for a given parameter set. The high-temperature scan is shown in Fig. 3, and the low-temperature scan in Fig. 4. To illustrate the dependence on key parameters, we color-code the points by the ratio $T_{p}/m_{A^{\prime}}$ in the left panels and by $v_{x}$ in the right panels of Figs. 3 and 4. We recover the same qualitative trend observed in the gauge-dependent scan: larger $v_{x}$ typically shifts the signal toward higher peak frequencies.
It is important to note, however, that the raw outputs of these scans include parameter points that lie outside the temperature-limit assumptions underlying the corresponding gauge-independent treatments. In particular, the high-temperature scan contains points with $T_{p}/m_{A^{\prime}}<2$, which do not satisfy the (approximately) high-temperature criterion and must be discarded. Likewise, the low-temperature scan contains points with $T_{p}/m_{A^{\prime}}>0.5$, which fall outside the (approximately) low-temperature domain and must be discarded.
To ensure consistency with the assumptions adopted in the preceding analysis, we retain only points with $T_{p}/m_{A^{\prime}}\gtrsim 2$ as belonging to the (approximately) high-temperature regime, and points with $T_{p}/m_{A^{\prime}}\lesssim 0.5$ as belonging to the (approximately) low-temperature regime. Within the latter, points with $T_{p}/m_{A^{\prime}}\lesssim 0.2$ correspond to genuinely supercooled transitions. The surviving points from the two scans are combined in Fig. 5, which provides concrete, gauge-independent predictions for the minimal gauged $U(1)_{x}$ dark sector.
The empty gap in Fig. 5 corresponds to the intermediate regime $0.5<T_{p}/m_{A^{\prime}}<2$, where the high-temperature expansion is not valid and the low-temperature thermal contributions cannot be treated as small corrections. Consequently, neither the high- nor low-temperature gauge-independent framework applies, and concrete predictions in this intermediate regime are not currently under theoretical control. Moreover, the corresponding gravitational wave signals are typically weaker than those from supercooled transitions and are generally below the projected reach of near-future experiments, rendering this regime less phenomenologically relevant at present.
Figure 3: Gauge-independent gravitational wave signals in the high-temperature regime. Each scatter point shows the peak frequency and peak amplitude of the stochastic background sourced by a first-order phase transition. In panel (a), points are color-coded by $T_{p}/m_{A^{\prime}}$. In panel (b), points are color-coded by the $U(1)_{x}$ symmetry breaking scale $v_{x}$, scanned over $v_{x}\in[10^{-3},10^{2}]\penalty 10000\ \mathrm{GeV}$. All points with $T_{p}/m_{A^{\prime}}<2$ violate the high-temperature assumption and should be discarded.
Figure 4: Gauge-independent gravitational wave signals in the low-temperature regime. Each scatter point shows the peak frequency and peak amplitude of the stochastic background sourced by a first-order phase transition. In panel (a), points are color-coded by $T_{p}/m_{A^{\prime}}$. In panel (b), points are color-coded by the $U(1)_{x}$ symmetry breaking scale $v_{x}$, scanned over $v_{x}\in[10^{-3},10^{2}]\penalty 10000\ \mathrm{GeV}$. All points with $T_{p}/m_{A^{\prime}}>0.5$ violate the low-temperature assumption and should be discarded.
Figure 5: Concrete gauge-independent gravitational wave predictions for the minimal gauged $U(1)_{x}$ dark sector, obtained by combining the physically consistent points selected from Fig. 3 ($T_{p}/m_{A^{\prime}}\gtrsim 2$) and Fig. 4 ($T_{p}/m_{A^{\prime}}\lesssim 0.5$). The gap corresponds to the intermediate regime $0.5<T_{p}/m_{A^{\prime}}<2$, where neither the high- nor low-temperature gauge-independent treatment is applicable. Low-temperature points are shown in blue, with darker shades corresponding to smaller $T_{p}/m_{A^{\prime}}$ (stronger supercooling). High-temperature points are shown in red, with darker shades corresponding to larger $T_{p}/m_{A^{\prime}}$. Black crosses mark three benchmark models for Case 2, for which the dark matter evolution is computed explicitly (see Table 3), providing benchmark targets for multi-messenger searches of the corresponding $U(1)$ model. In particular, benchmark models (e) and (f) lie within the projected sensitivity of Taiji, TianQin, and LISA.
To illustrate the impact of gauge dependence on the predicted gravitational wave signals, we explicitly compare our gauge-independent treatment with the gauge-dependent calculation in Fig. 6. As representative examples, we consider benchmark models (e) and (f), which undergo supercooled phase transitions, and present the resulting gravitational wave spectra $\Omega_{\rm GW}h^{2}$ as functions of frequency. The gauge-independent predictions are shown as solid curves. The gauge-dependent results are obtained by varying the gauge parameter over $\xi\in[0,5]$ and are displayed as shaded bands; the dashed curves indicate the Landau gauge case with $\xi=0$. One observes substantial deviations between the gauge-independent predictions and the gauge-dependent results.
Figure 6: Comparison of the gauge-independent treatment with the gauge-dependent computation. Shown are the gravitational wave spectra $\Omega_{\rm GW}h^{2}$ as functions of frequency for benchmark models (e) and (f) in the low-temperature regime. The shaded bands indicate the gauge-dependent results obtained by varying the gauge parameter over $\xi\in[0,5]$. Solid curves denote the gauge-independent predictions, while dashed curves show the gauge-dependent results in Landau gauge ($\xi=0$).
5 Evolution of dark sector particles
A successful dark sector model should also furnish viable dark matter candidates. With the steadily improving sensitivity of dark matter detection experiments and collider searches, many conventional freeze-out scenarios, especially in the traditional WIMP mass range $\mathcal{O}(10-100\penalty 10000\ \mathrm{GeV})$, are increasingly constrained. In large regions of parameter space, the annihilation of dark matter into SM states,
$XX\penalty 10000\ ({\rm or}\penalty 10000\ X\overline{X})\xrightarrow{\ \text{annihilation}\ }\text{SM particles}\,$, (5.1)
is no longer efficient enough to yield the observed relic abundance without conflicting with existing bounds [109, 110].
Within the minimal $U(1)_{x}$ setup, two well-motivated possibilities arise, as discussed in Section 2. First, the dark photon can serve as dark matter provided it is stable, which typically requires vanishing or negligibly small kinetic mixing with the hypercharge gauge field. In this limit its relic abundance is set by Higgs portal interactions. Second, a vectorlike dark fermion can be the dark matter candidate, communicating with the SM predominantly through the dark photon portal, since it does not couple directly to the dark Higgs in the minimal field content, a non-negligible kinetic mixing is then required to maintain efficient interactions. Alternatively, one may introduce two vectorlike fermions $\chi_{1}$ and $\chi_{2}$ with an allowed Yukawa interaction $\sim H_{x}\,\overline{\chi_{1}}\chi_{2}$ for an appropriate assignment of $U(1)_{x}$ charges.
In this paper, we consider both classes of dark matter candidates in the secluded regime, where they annihilate into light dark sector mediators that subsequently decay into SM states.
5.1 Boltzmann equations for dark particles
5.1.1 Dark photon dark matter from the Higgs portal
The dominant annihilation channel for dark photon dark matter proceeds via
$A^{\prime}A^{\prime}\xrightarrow{\ \text{annihilation}\ }h_{x}h_{x}\xrightarrow{\quad\text{decay}\quad}\text{SM\ particles}\,.$ (5.2)
The coupled Boltzmann equations governing $Y_{h_{x}}$ and $Y_{A^{\prime}}$ are given by
$\frac{\mathrm{d}Y_{h_{x}}}{\mathrm{d}T}=$ $\;-\frac{s}{T\overline{H}}\sum_{i\in\mathrm{SM}}\Bigl\{2\bigl[(Y_{h_{x}}^{\mathrm{eq}})^{2}-Y_{h_{x}}^{2}\bigr]\left\langle\sigma v\right\rangle_{h_{x}h_{x}\to\{i\overline{i},hh,ZZ,W^{\pm}W^{\mp}\}}$
$+(Y_{h}Y_{h_{x}}^{\mathrm{eq}}-Y_{h}Y_{h_{x}})\left\langle\sigma v\right\rangle_{hh_{x}\to\{i\overline{i},hh,ZZ,W^{\pm}W^{\mp}\}}$
$+2Y_{A^{\prime}}^{2}\left\langle\sigma v\right\rangle_{A^{\prime}A^{\prime}\to h_{x}h_{x}}-2Y_{h_{x}}^{2}\left\langle\sigma v\right\rangle_{h_{x}h_{x}\to A^{\prime}A^{\prime}}$
$+Y_{A^{\prime}}^{2}\left\langle\sigma v\right\rangle_{A^{\prime}A^{\prime}\to hh_{x}}-Y_{h}Y_{h_{x}}\left\langle\sigma v\right\rangle_{hh_{x}\to A^{\prime}A^{\prime}}$
$+2\,\theta(m_{h}-2m_{h_{x}})\bigl(-Y_{h_{x}}^{2}\left\langle\sigma v\right\rangle_{h_{x}h_{x}\to h}+\tfrac{1}{s}Y_{h}\langle\Gamma\rangle_{h\to h_{x}h_{x}}\bigr)$
$+\theta(m_{h_{x}}-2m_{i})\,\tfrac{1}{s}(Y_{h_{x}}^{{\rm eq}}-Y_{h_{x}})\langle\Gamma\rangle_{h_{x}\to i\overline{i}}$
$+\theta(m_{h_{x}}-2m_{A^{\prime}})\bigl(Y_{A^{\prime}}^{2}\left\langle\sigma v\right\rangle_{A^{\prime}A^{\prime}\to h_{x}}-\tfrac{1}{s}Y_{h_{x}}\langle\Gamma\rangle_{h_{x}\to A^{\prime}A^{\prime}}\bigr)\Bigr\}\,$, (5.3)
$\frac{\mathrm{d}Y_{A^{\prime}}}{\mathrm{d}T}=$ $\;-2\times\frac{s}{T\overline{H}}\sum_{i\in\mathrm{SM}}\Bigl\{\bigl[(Y_{A^{\prime}}^{\mathrm{eq}})^{2}-Y_{A^{\prime}}^{2}\bigr]\left\langle\sigma v\right\rangle_{A^{\prime}A^{\prime}\to\{i\overline{i},hh,ZZ,W^{\pm}W^{\mp}\}}$
$-Y_{A^{\prime}}^{2}\left\langle\sigma v\right\rangle_{A^{\prime}A^{\prime}\to h_{x}h_{x}}+Y_{h_{x}}^{2}\left\langle\sigma v\right\rangle_{h_{x}h_{x}\to A^{\prime}A^{\prime}}$
$-Y_{A^{\prime}}^{2}\left\langle\sigma v\right\rangle_{A^{\prime}A^{\prime}\to hh_{x}}+Y_{h}Y_{h_{x}}\left\langle\sigma v\right\rangle_{hh_{x}\to A^{\prime}A^{\prime}}$
$+\theta(m_{h}-2m_{A^{\prime}})\bigl(-Y_{A^{\prime}}^{2}\left\langle\sigma v\right\rangle_{A^{\prime}A^{\prime}\to h}+\tfrac{1}{s}Y_{h}\langle\Gamma\rangle_{h\to A^{\prime}A^{\prime}}\bigr)$
$+\theta(m_{h_{x}}-2m_{A^{\prime}})\bigl(-Y_{A^{\prime}}^{2}\left\langle\sigma v\right\rangle_{A^{\prime}A^{\prime}\to h_{x}}+\tfrac{1}{s}Y_{h_{x}}\langle\Gamma\rangle_{h_{x}\to A^{\prime}A^{\prime}}\bigr)\Bigr\}\,$, (5.4)
where
$\overline{H}=\frac{H}{1+\frac{1}{3}\frac{T}{h_{{\rm eff}}}\frac{{\rm d}h_{{\rm eff}}}{{\rm d}T}}=\sqrt{\frac{\pi^{2}g_{{\rm eff}}}{90}}\frac{T^{2}/M_{{\rm Pl}}}{1+\frac{1}{3}\frac{T}{h_{{\rm eff}}}\frac{{\rm d}h_{{\rm eff}}}{{\rm d}T}}\,.$ (5.5)
5.1.2 Dark fermion dark matter from the $U(1)$ portal
The benchmark models we consider are chosen such that the dark Higgs $h_{x}$ remains in thermal equilibrium and decays into SM fermions before BBN, and thus without affecting the number density evolution of the dark matter or the dark photon. We therefore focus on the coupled evolution of $\chi$ and $A^{\prime}$. The dominant annihilation channel for dark fermion dark matter is
$\chi\bar{\chi}\xrightarrow{\quad Z,\ A^{\prime}\quad}{\rm SM\ particles}\,.$ (5.6)
The coupled Boltzmann equations for $Y_{\chi}$ and $Y_{A^{\prime}}$ are given by
$\frac{\mathrm{d}Y_{\chi}}{\mathrm{d}T}=$ $\;-\frac{s}{T\overline{H}}\sum_{i\in{\rm SM}}\Bigl\{\bigl[(Y_{\chi}^{\mathrm{eq}})^{2}-Y_{\chi}^{2}\bigr]\langle\sigma v\rangle_{\chi\overline{\chi}\rightarrow\{i\overline{i},ZZ\}}$
$-Y_{\chi}^{2}\langle\sigma v\rangle_{\chi\overline{\chi}\to A^{\prime}Z}+Y_{A^{\prime}}Y_{Z}\langle\sigma v\rangle_{A^{\prime}Z\to\chi\overline{\chi}}$
$-Y_{\chi}^{2}\langle\sigma v\rangle_{\chi\overline{\chi}\to A^{\prime}A^{\prime}}+Y_{A^{\prime}}^{2}\langle\sigma v\rangle_{A^{\prime}A^{\prime}\to\chi\overline{\chi}}$
$+\theta(m_{A^{\prime}}-2m_{\chi})\bigl(-Y_{\chi}^{2}\langle\sigma v\rangle_{\overline{\chi}\chi\rightarrow A^{\prime}}+\tfrac{1}{s}Y_{A^{\prime}}\langle\Gamma\rangle_{A^{\prime}\to\chi\overline{\chi}}\bigr)\Big\}\,$, (5.7)
$\frac{\mathrm{d}Y_{A^{\prime}}}{\mathrm{d}T}=$ $\;-\frac{s}{T\overline{H}}\sum_{i\in{\rm SM}}\Bigl\{(Y_{A^{\prime}}^{{\rm eq}}-Y_{A^{\prime}})\left[Y_{Z}\langle\sigma v\rangle_{A^{\prime}Z\to i\overline{i}}+Y_{\gamma}\langle\sigma v\rangle_{A^{\prime}\gamma\to i\overline{i}}\right$.
$\left.+2Y_{i}\bigl(\langle\sigma v\rangle_{iA^{\prime}\to iZ}+\langle\sigma v\rangle_{iA^{\prime}\to i\gamma}\bigr)\right]$
$+Y_{\chi}^{2}\langle\sigma v\rangle_{\chi\overline{\chi}\to A^{\prime}A^{\prime}}-Y_{A^{\prime}}^{2}\langle\sigma v\rangle_{A^{\prime}A^{\prime}\to\chi\overline{\chi}}$
$+\theta(m_{A^{\prime}}-2m_{\chi})\bigl(Y_{\chi}^{2}\langle\sigma v\rangle_{\overline{\chi}\chi\rightarrow A^{\prime}}-\tfrac{1}{s}Y_{A^{\prime}}\langle\Gamma\rangle_{A^{\prime}\to\chi\overline{\chi}}\bigr)$
$+\theta(m_{A^{\prime}}-2m_{i})\bigl(Y_{i}^{2}\langle\sigma v\rangle_{i\overline{i}\rightarrow A^{\prime}}-\tfrac{1}{s}Y_{A^{\prime}}\langle\Gamma\rangle_{A^{\prime}\to i\overline{i}}\bigr)\Big\}\,.$ (5.8)
5.2 Benchmark models
Case-1: dark photon dark matter
Current bounds on the mixing parameter $|\sin\theta\,|$, c.f., Eq. (2.25) arise from rare meson decays, beam-dump experiments, collider searches, and astrophysical observations, see [111, 112, 113] for recent summaries. Colliders also constrain invisible Higgs decays. The ATLAS Collaboration reports ${\rm Br}(h\to{\rm inv})<0.107$ at $95\%$ CL [114], while CMS finds ${\rm Br}(h\to{\rm inv})<0.15$ at $95\%$ CL [115]. The invisible branching fraction is
${\rm Br}(h\to{\rm inv})=\frac{\Gamma_{h\to{\rm inv}}}{\Gamma_{h\to{\rm SM}}+\Gamma_{h\to{\rm inv}}}\,,\qquad\Gamma_{h\to{\rm inv}}=\Gamma_{h\to h_{x}h_{x}}+\Gamma_{h\to A^{\prime}A^{\prime}}\,$, (5.9)
where for $m_{h}\simeq 125\penalty 10000\ {\rm GeV}$ the SM Higgs width is $\Gamma_{h\to{\rm SM}}=4.1\penalty 10000\ {\rm MeV}$ [116]. Since the dark Higgs $h_{x}$ ultimately decays into SM fermions, its lifetime must be shorter than $\mathcal{O}(1\penalty 10000\ {\rm s})$ to avoid spoiling big-bang nucleosynthesis.
For this scenario we present three benchmark models in Table 2, with $v_{x}=$ 100, 200, 450 GeV and a mass hierarchy $m_{A^{\prime}}>m_{h_{x}}$. This hierarchy is required to realize a first-order phase transition and hence to generate a stochastic gravitational wave signal. All benchmarks are consistent with current experimental bounds.
For each benchmark we compute both the number density evolution of the relevant dark sector species and the gravitational wave signal from the phase transition. The Boltzmann equations governing $Y_{h_{x}}$ and $Y_{A^{\prime}}$ are given in Eqs. (5.3)–(5.4). As the Universe cools, the dark sector undergoes a first-order phase transition during which $U(1)_{x}$ is spontaneously broken and the dark photon $A^{\prime}$ acquires a mass. After the transition, $A^{\prime}$ annihilates into lighter states, including the dark Higgs $h_{x}$ and SM particles, and subsequently freezes out, yielding the observed dark matter relic abundance. The dark Higgs $h_{x}$ is unstable and decays into SM fermions before big bang nucleosynthesis (BBN).
Although these benchmarks satisfy the dark matter relic density constraint, they typically require a tiny dark gauge coupling $g_{x}$ to avoid excessive secluded depletion of the the dark photon. As a result, the phase transition is relatively weak and the corresponding gravitational wave signal is suppressed. Moreover, for these benchmarks we find that $T_{p}/m_{A^{\prime}}$ lies outside the validity ranges of our gauge-independent treatments in both the high- and low-temperature limits.
Model $v_{x}$ $g_{x}$ $\lambda_{x}$ $\lambda_{{\rm mix}}$ $m_{h_{x}}$ $m_{A^{\prime}}$ $\Omega_{A^{\prime}}h^{2}$ $\tau_{h_{x}}$ $T_{p}$ $T_{p}/m_{A^{\prime}}$
(a) $100$ $0.0221$ $84\times 10^{-7} 1×1$$\times 10^{-5}$ $0.061$ $2.21$ $0.12$ $2.61\times 10^{-3}$ $3.90$ $1.77$
(b) $200$ $0.0372$ $00\times 10^{-6} 1×1$$\times 10^{-5}$ $0.28$ $7.44$ $0.12$ $5.19\times 10^{-7}$ $10.85$ $1.46$
(c) $450$ $0.0525$ $2.95\times 10^{-6}$ $1\times 10^{-6}$ $1.09$ $23.63$ $0.12$ $9.64\times 10^{-7}$ $29.69$ $1.26$
Table 2: Benchmark models for Case 1, in which the dark photon is the dark matter candidate and the dark sector communicates with the SM through the Higgs portal. All masses are in GeV. For each benchmark we compute both the dark sector number density evolution and the first-order phase transition dynamics. The Higgs mixing parameter $\lambda_{\rm mix}$ satisfies current experimental bounds, and the $h_{x}$ lifetimes (in seconds) are ensured to be shorter than $1\penalty 10000\ \mathrm{s}$. We also list the percolation temperature $T_{p}$ and the ratio $T_{p}/m_{A^{\prime}}$ for each benchmark.
Case-2: dark fermion dark matter
In Case 2, the dark matter candidate is a vectorlike dark fermion $\chi$ with no direct Yukawa coupling to the dark Higgs $h_{x}$. As a result, $\chi$ affects the finite-temperature effective potential $V_{\rm eff}(\phi_{c},T)$ only indirectly through its contribution to the longitudinal dark photon thermal mass $\Pi_{A^{\prime}}$, c.f., Eq. (3.15). Consequently, in the gauge-independent calculation the coefficient appearing in $\Pi_{A^{\prime}}$ is modified to $2/3$.
The kinetic-mixing parameter $\delta$ in this work denotes the mixing between $U(1)_{x}$ and the hypercharge gauge field $U(1)_{Y}$, c.f., Eq. (2.4). Experimental searches constrain the kinetic mixing between the dark photon and the electromagnetic photon, parameterized by $\bm{\delta}$. The two parameters are related by [66, 117]
$\delta\approx\frac{\sqrt{g_{2}^{2}+g_{Y}^{2}}}{g_{2}}\bm{\delta}\,.$ (5.10)
For Case 2, we choose three benchmark models, listed in Table 3, with $v_{x}=$ 1, 10, 100 GeV. For each benchmark we compute both the dark sector number density evolution and the associated gravitational wave signal. All benchmarks satisfy current experimental constraints on the kinetic mixing parameter $\delta$ and are shown in Fig. 7. The Boltzmann equations governing the evolution of $Y_{\chi}$ and $Y_{A^{\prime}}$ are given in Eqs. (5.7)–(5.8).
In model (d), the dark fermion is heavier than the dark photon, $m_{\chi}>m_{A^{\prime}}$. As the Universe cools, $\chi\bar{\chi}$ annihilates efficiently into dark photons and yields the observed relic abundance. The dark photons subsequently decay into $e^{+}e^{-}$ pairs before BBN. For models (e) and (f), we instead choose $m_{\chi}$ to lie slightly below $m_{A^{\prime}}$. This near-degeneracy allows for a larger $g_{x}$ while still reproducing the correct $\Omega_{\chi}h^{2}$ via the process $A^{\prime}A^{\prime}\to\chi\overline{\chi}$. We indicate these three Case 2 benchmarks in Figs. 5 and 7. Benchmark models (e) and (f) lie within the projected sensitivity of Taiji, TianQin, and LISA, and thus provide promising multi-messenger targets for a genuinely minimal $U(1)$ dark sector.
Model $v_{x}$ $g_{x}$ $\lambda_{x}$ $\delta$ $m_{\chi}$ $m_{A^{\prime}}$ $\Omega_{\chi}h^{2}$ $\tau_{A^{\prime}}$ $T_{p}$ $T_{p}/m_{A^{\prime}}$
(d) 11 $0.0208$ $1.93\times 10^{-7}$ $1.3\times 10^{-10}$ $0.35$ $0.0208$ $0.12$ $0.35$ $0.0423$ $2.03$
(e) $10$ 0.505 $17\times 10^{-4} 1×1$$\times 10^{-8}$ $4.50$ 5.05 0.12 $1.05\times 10^{-7}$ $0.54$ $0.107$
(f) $100$ 0.391 $3.13\times 10^{-5}$ $1\times 10^{-8}$ $38.50$ 39.10 0.12 $1.14\times 10^{-8}$ $3.34$ $0.085$
Table 3: Benchmark models for Case 2, in which the dark matter candidate is a vectorlike fermion $\chi$ that communicates with the SM through $U(1)$ portal via kinetic mixing. All masses are in GeV and lifetimes in seconds. For each benchmark we compute both the dark sector number density evolution and the first-order phase transition dynamics. All benchmarks satisfy the relevant constraints on the kinetic mixing parameter $\delta$ (see Fig. 7). The corresponding peak gravitational wave signals are shown in Fig. 5. We also list the percolation temperature $T_{p}$ and the ratio $T_{p}/m_{A^{\prime}}$ for each benchmark.
Figure 7: Experimental constraints on the kinetic mixing parameter $\delta$, defined in Eqs. (2.4) and (5.10), for dark photon masses $m_{A^{\prime}}\in[1{\rm MeV},\,100{\rm GeV}]$. The three Case 2 benchmark models (Table 3) are shown as black crosses and are consistent with all current experimental bounds. Constraints on the kinetic mixing for dark photons with $m_{A^{\prime}}\gtrsim$ 1 MeV arise from supernova (SN1987A) [118], $(g-2)_{e}$ [119] and from various beam dump experiments: E141 [120], E137 [121, 122, 123], $\nu$-Cal [124, 125]. The pink region in the upper portion of the plot, labeled as “Others”, collects constraints from collider and fixed-target experiments, including: CMS [126], BaBar [127], LHCb [128], KLOE [129, 130, 131, 132] and A1 [133].
5.3 Phenomenological implications
The continuing non-observation of weakly interacting massive particles (WIMP), especially in the $\mathcal{O}(10-100\penalty 10000\ {\rm GeV})$ region, has substantially challenged a large class of traditional dark matter scenarios. Current limits from direct and indirect detection, together with collider searches, have significantly narrowed the viable parameter space for many WIMP candidates. As a result, the measured relic abundance has increasingly become a stringent constraint on model building rather than an automatic target that can be achieved across broad regions of parameter space.
At the level of thermal freeze-out, the tension is particularly transparent: with present experimental sensitivities, many WIMP-like candidates that remain compatible with direct detection bounds cannot maintain sufficient annihilation into SM final states, which leads to an overproduction of dark matter. In hidden sector constructions, this difficulty can be even more pronounced, since the interactions connecting the dark sector to the SM are typically mediated by portal couplings that must be small. Consequently, annihilation directly into SM particles can be inefficient. An common alternative is secluded annihilation, in which dark matter annihilates dominantly into lighter dark mediators that subsequently decay into SM states, thereby achieving an efficient depletion of the relic abundance while alleviating dark matter experimental constraints.
In the minimal $U(1)_{x}$ framework considered here, these general considerations manifest themselves clearly in the benchmark models. As illustrated in Fig. 5, parameter points with dark fermion dark matter can simultaneously achieve sufficient annihilation to match the observed relic density and produce a potentially detectable stochastic gravitational wave signal from the first-order phase transition. By contrast, dark photon dark matter realized through secluded freeze-out cannot accommodate both requirements: regions of parameter space that yield an observable gravitational wave signal often require a sizable gauge coupling, while such a coupling tends to deplete the dark photon relic abundance too efficiently, making it difficult to retain the correct relic density. This complementarity is not surprising. Detectable gravitational wave signals from a first-order phase transition generally favor a sizable gauge coupling, which strengthens the dynamics of the transition and enhances the signal. However, the same sizable coupling often increases the efficiency of dark photon annihilation, reducing the relic density below the observed value via freeze-out.
Dark fermion dark matter can annihilate into dark photons, which then decay into SM particles via portal interactions. The presence of an additional mass parameter $m_{\chi}$ offers additional flexibility. In particular, choosing $m_{\chi}$ close to the dark photon mass can suppress the annihilation efficiency and prevent excessive depletion, enabling simultaneous compatibility with the relic abundance and gravitational wave detectability.
It should be noted that in the minimal $U(1)_{x}$ setup, the dark photon and dark Higgs masses are $m_{A^{\prime}}=g_{x}v_{x}$ and $m_{h_{x}}=\sqrt{2\lambda}\,v_{x}$, respectively. Given the scaling relation between the Higgs quartic coupling and the gauge coupling discussed in detail in Section 3 in both high- and low-temperature regimes, the dark photon is generically heavier than the dark Higgs. Accordingly, in all benchmark models considered here we impose $m_{A^{\prime}}>m_{h_{x}}$, a hierarchy that is required for a dark sector phase transition in this minimal setup.
Finally, we emphasize that production via freeze-in can remain viable for both classes of dark matter candidates even when the gauge coupling is sizable. For dark fermion dark matter, previous studies, see, e.g., [134, 66], [^9] Although these references consider a simplified $U(1)_{x}$ setup consisting only of a vectorlike fermion and a dark photon whose mass arises from the Stückelberg mechanism, they already illustrate the broad viability and flexibility of freeze-in production in $U(1)$ portal scenarios. demonstrate that large regions of parameter space can satisfy existing experimental constraints while reproducing the observed relic abundance. For dark photon dark matter, even if the hidden sector possesses a sizable gauge coupling, the Higgs portal mixing can be adjusted to generate an appropriate abundance. More broadly, freeze-in production can offer greater freedom to maintain a sizable gauge coupling, and hence potentially large gravitational wave signals, while still achieving the correct relic density. A systematic exploration of this possibility, in particular for dark photon dark matter in the minimal gauged $U(1)_{x}$ setup, is left open for future work.
6 Conclusion and outlook
In this work, we investigated gauge-independent predictions for stochastic gravitational waves sourced by first-order phase transitions in a minimal gauged $U(1)_{x}$ dark sector. This addresses a central obstacle to making reliable predictions in gauge theories: the gauge dependence of the finite-temperature effective potential and the associated tunneling action. To obtain model-intrinsic and reproducible results, we adopted a gauge-independent formulation of the effective action based on the Nielsen identity, combined with a controlled derivative expansion and power counting. This framework yields a gauge-independent tunneling action, enabling a well-defined treatment of nucleation dynamics and, consequently, robust gravitational wave predictions. The resulting signals span a wide range of peak frequencies relevant to pulsar timing arrays and planned space-based interferometers, and they exhibit systematic trends across parameter space.
We performed Monte Carlo explorations of the parameter space using both the conventional gauge-dependent approach and, separately, the gauge-independent framework in the high- and low-temperature limits. The gauge-dependent scan is useful for building intuition and reveals clear qualitative trends: the $U(1)_{x}$ symmetry breaking scale $v_{x}$ largely controls the characteristic frequency, while the phase transition temperature regime correlates strongly with the signal strength, with supercooled transitions typically producing larger amplitudes. At the same time, an explicit comparison shows that gauge-dependent spectra can depart significantly from the gauge-independent predictions when the gauge-fixing parameter is varied, underscoring that the gauge choice can materially impact phenomenological interpretations.
Using the gauge-independent effective action, we scanned the high- and low-temperature limits separately and then retained only those points consistent with the corresponding temperature-limit assumptions. The surviving points shown in Fig. 5 provide concrete, gauge-independent predictions for the minimal $U(1)_{x}$ setup, mapping viable microscopic parameters to detector-facing peak frequencies and amplitudes. A notable outcome is that supercooled transitions constitute a substantial fraction of successful transitions in broad scans and, when realized, often populate the most promising regions of experimental sensitivity, reinforcing the view that supercooled first-order phase transitions are among the most compelling targets in current gravitational wave searches.
By contrast, parametrically high-temperature transitions are comparatively rare and typically yield weak signals. Between these limits lies an intermediate regime in which neither the high- nor the low-temperature gauge-independent treatment is applicable. Bridging this gap with a fully gauge-independent framework remains an important open direction.
Finally, we connected the phase transition phenomenology to viable dark matter candidates within the same minimal field content. We considered both dark photon dark matter via the Higgs portal, and dark fermion dark matter via the dark photon portal, in secluded settings, and provided benchmark targets that enable genuine multi-messenger tests. The benchmarks illustrate a clear complementarity: parameter regions capable of producing potentially detectable gravitational waves typically favor sizable gauge dynamics, which can be in tension with reproducing the observed relic abundance for dark photon dark matter via secluded freeze-out. By contrast, dark fermion dark matter retains greater flexibility and can more readily accommodate both relic density consistency and gravitational wave visibility. More broadly, freeze-in production channels may further expand the viable parameter space while maintaining sizable gauge couplings, and merit systematic exploration in future work.
Across all temperature regimes, we find a robust feature of the parameter region that supports a first-order phase transition: the dark photon must be heavier than the dark Higgs. If nature were to realize the opposite hierarchy, a $U(1)_{x}$ breaking phase transition in the minimal dark sector and thus the associated gravitational wave signal, will not occur.
Taken together, this work delivers a gauge-independent, end-to-end pipeline that connects a minimal hidden sector Lagrangian to nucleation dynamics, stochastic gravitational wave spectra, and cosmologically viable dark matter benchmarks. Our results make clear that gauge independence is not merely a technical refinement, but a necessary condition for obtaining physically meaningful gravitational wave predictions in gauge theories. Within this framework, even the minimal gauged $U(1)_{x}$ dark sector yields predictive, detector-relevant targets, most notably in the supercooled regime, that are complementary to collider probes and conventional dark matter searches.
Looking ahead, the setup studied here admits several natural extensions, including additional hidden states, gauge-independent conformal $U(1)$ realizations, and explorations of phase transitions at larger symmetry breaking scales. More broadly, the lesson is general: whenever gauge fields and spontaneous symmetry breaking play a dynamical role, gravitational wave predictions should be formulated in gauge-independent terms to enable reliable comparisons with experimental data and across different theoretical studies.
Acknowledgments:
This work is supported in part by the National Natural Science Foundation of China under Grant No. 11935009, and Tianjin University Self-Innovation Fund Extreme Basic Research Project Grant No. 2025XJ21-0007.